### v1

# Universal 3-Edge-Weightability of Regular Graphs

Abstract.
A graph $G=(V,E)$ is universally $3$-edge-weightable if for every set $\{a,b,c\}$ of three distinct real numbers, there exists an edge-weighting $w:E\to\{a,b,c\}$ such that $d_{w}(u)\neq d_{w}(v)$ for all $uv\in E$. The 1-2-3 Conjecture, recently settled by Keusch, ensures that all nice graphs admit a proper $\{1,2,3\}$-edge-weighting. For regular graphs, this implies proper edge-weightings for arithmetic progressions $\{a,b,c\}$ via affine transformation, but not for general triples due to the spacing-ratio preservation property of affine maps. We prove that every $k$-regular graph with $k\geq 2$ is universally $3$-edge-weightable. This establishes proper edge-weightings for all triples $\{a,b,c\}$, not just arithmetic progressions, showing that regular graphs satisfy a strictly stronger weighting property than what follows from the 1-2-3 Conjecture alone. The proof adapts the independent-set decomposition framework to edge-weightings and introduces a two-phase conflict-resolution scheme that operates independently of the specific choice of $\{a,b,c\}$.
Key words and phrases:
edge weighting, regular graph, 1-2-3 conjecture, universal 3-edge-weightability
2020 Mathematics Subject Classification:
05C15,05C78
Kecai Deng is the corresponding author.
  Kecai Deng is partially supported by NSFC (No. 11701195), and Fundamental Research Funds for the Central Universities (No. ZQN-904).
1. Introduction
All graphs considered are simple, finite, and undirected. A graph is nice if none of its components is isomorphic to $K_{2}$. For a graph $G=(V,E)$ and a subset $Q\subset\mathbb{R}$, a mapping $w:E\to Q$ is called a $Q$-edge weighting of $G$; the weighted degree of a vertex $v$ is defined as $d_{w}(v)=\sum_{e\ni v}w(e)$; an edge $e=uv$ is a conflict if $d_{w}(u)=d_{w}(v)$. If there is no conflict under $w$, then $w$ is called proper. We say $G$ is $Q$-edge-weightable if it admits a proper $Q$-edge weighting, and call $G$ universally $k$-edge-weightable if it is $Q$-edge weightable for every $k$-element subset $Q\subset\mathbb{R}$.
The 1-2-3 Conjecture, posed by Karoński, Łuczak, and Thomason [5], asserted that every nice graph is $\{1,2,3\}$-edge-weightable. After remaining open for two decades, this conjecture was finally confirmed by Keusch [6] in 2024. The property of being $\{1,2,3\}$-edge-weightable corresponds precisely to universal 3-edge-weightability restricted to the specific set $\{1,2,3\}$. For regular graphs, the 1-2-3 result can be extended via an affine transformation to show that they are $\{a,b,c\}$-edge-weightable for any arithmetic progression $\{a,b,c\}$ (where ba=cbb-a=c-b). However, this approach fails for general (non-arithmetic) three-element sets, leaving open the question of whether regular graphs are universally 3-edge-weightable.
Various strengthenings of the 1-2-3 Conjecture have been investigated. A graph is edge-weight $k$-choosable if for any assignment $L$ that gives each edge $e$ a list $L(e)$ of $k$ real numbers, there exists a proper edge weighting $w$ with $w(e)\in L(e)$ for all $e\in E$. The list version of the 1-2-3 Conjecture [1] asks whether every nice graph is edge-weight $3$-choosable. A stronger concept is $(1,3)$-choosability [7]: a graph is $(1,3)$-choosable if, for an arbitrary vertex weighting $f:V\rightarrow\mathbb{R}$ and any assignment of $3$-element lists to edges, there exists an edge weighting $w$ using weights from the respective lists such that $f(u)+d_{w}(u)\neq f(v)+d_{w}(v)$ for every edge $uv\in E$.
These properties form a clear hierarchy:
$(1,3)\text{-choosable}\ \Rightarrow\ 3\text{-choosable}\ \Rightarrow\ \text{universally 3-edge-weightable}\ \Rightarrow\ \{1,2,3\}\text{-edge-weightable}.$
Progress on the strongest property, $(1,3)$-choosability, has been steady. Cao [3] proved that every nice graph is $(1,17)$-choosable, improved by Zhu [8] to $(1,5)$-choosability. Consequently, every nice graph is edge-weight $5$-choosable. However, it remains open whether every nice graph is $(1,3)$-choosable, or even edge-weight $3$-choosable. Partial results show that several graph classes are $(1,3)$-choosable, including complete graphs and complete bipartite graphs [1], generalized Halin graphs [12], graphs with maximum average degree less than $\frac{11}{4}$ [9], dense Eulerian graphs [13], and Cartesian products of paths and cycles [11].
Between the weakest property (settled $\{1,2,3\}$-edge-weightability) and the stronger open problem of $3$-choosability lies the natural question of universal 3-edge-weightability:
Problem 1.1 (Universal 3-edge-weightability).
Is every nice graph universally 3-edge-weightable?
In other words, can the fixed set $\{1,2,3\}$ be replaced by an arbitrary 3-element set? Our result (Corollary 2) shows that regular graphs are universally 3-edge-weightable, while Problem 1 asks whether this holds for all nice graphs.
A related but different line of research concerns total weightings, where weights are assigned to both vertices and edges. For this total weighting problem, the 1-2 Conjecture [10] asks whether every nice graph is $\{1,2\}$-total-weightable. Notably, for regular graphs a stronger universal 2-total-weightability has been established [4], which serves as motivation for the present work on universal 3-edge-weightability.
Our contribution
In this paper, we demonstrate that the independent set decomposition framework in [4] can be adapted from total weighting to edge weighting, and can furthermore handle universal 3-edge-weighting. Our main technical result is the following:
Theorem 1.
For any positive real numbers $d_{1},d_{2}$ with $0<d_{1}<d_{2}$, every $k$-regular graph with $k\geq 3$ is $\{-d_{1},0,d_{2}\}$-edge-weightable.
A simple scaling and shifting argument yields the following corollary from Theorem 1.
Corollary 2.
Every $k$-regular graph ($k\geq 2$) is universally 3-edge-weightable.
Proof.
For $k=2$, a $2$-regular graph consists of disjoint cycles. Graphs with maximum average degree less than $\frac{11}{4}$ are known to be $(1,3)$-choosable [9]. Taking the vertex weighting $f\equiv 0$ and giving every edge the same list $\{a,b,c\}$, $(1,3)$-choosability implies the existence of a proper $\{a,b,c\}$-edge weighting. Thus we may assume $k\geq 3$.
It suffices to show that $G$ admits a proper $\{-(b-a),0,c-b\}$-edge weighting $w^{\prime}$. Indeed, if we then define $w=w^{\prime}+b$ (adding $b$ to every edge weight), the regularity of $G$ ensures that $w$ is a proper $\{a,b,c\}$-edge weighting.
Consider three cases. If ba<cbb-a<c-b, Theorem 1 directly provides a proper $\{-(b-a),0,c-b\}$-edge weighting. If ba=cbb-a=c-b, Keusch’s theorem [6] gives a proper $\{1,2,3\}$-edge weighting $w^{\prime\prime}$ of $G$. Setting $w^{\prime}=(b-a)(w^{\prime\prime}-2)$ yields a proper $\{-(b-a),0,c-b\}$-edge weighting. Finally, if ba>cbb-a>c-b, Theorem 1 supplies a proper $\{-(c-b),0,b-a\}$-edge weighting $w^{\prime\prime\prime}$; then $w^{\prime}=-w^{\prime\prime\prime}$ is the desired weighting. In all cases the required weighting exists, proving the corollary. ∎
The result of Corollary 2 places regular graphs strictly between the settled 1-2-3 Conjecture and the open $3$-choosability question in the hierarchy of edge-weighting properties.
Methodological remarks
Our proof of Theorem 1 adapts the independent-set decomposition framework from total-weightings to edge-weightings while addressing several distinctive challenges.
1. Absence of vertex-weight adjustments. In total weighting, vertex weights provide an independent degree of freedom that can be used to resolve conflicts. Edge weighting lacks this flexibility. To compensate, after partitioning the vertices into special independent sets $I_{0},I_{1},\ldots,I_{\ell}$, we employ the maximum independent set $I_{0}$ as a global regulator. We systematically weight $G[\cup_{i\geq 1}I_{i}]$ with $\{0,d_{2}\}$, and assign weight $d_{1}$ to selected edges through matchings between $I_{0}$ and $\cup_{i\geq 1}I_{i}$. This creates the separation $d_{w}(v)<d_{w}(u)$ for all $v\in I_{i}$ and $u\in I_{j}$ with $1\leq i<j$, and separation $d_{w}(x)<d_{w}(z)$ for all $v\in I_{0}$ and $z\in\cup_{i\geq 2}I_{i}$. The separation prevents conflicts between $I_{0}$ and $G[\cup_{i\geq 2}I_{i}]$, and conflicts in $G[\cup_{i\geq 1}I_{i}]$.
2. Different behavior under affine transformations. For total weighting on $k$-regular graphs, any proper $\{a_{0},b_{0}\}$-total weighting $w$ can be affinely transformed into a proper $\{a,b\}$-total weighting $w^{\prime}$. The transformation $w^{\prime}\mapsto\frac{b-a}{b_{0}-a_{0}}(w-a_{0})+a$, applied uniformly to vertex and edge weights, yields
$d_{w^{\prime}}(v)=\frac{b-a}{b_{0}-a_{0}}d_{w}(v)+(k+1)\left(a-\frac{b-a}{b_{0}-a_{0}}a_{0}\right),$
which preserves properness because the additive term is constant across vertices. Proving the $\{0,1\}$-case therefore suffices for the universal $\{a,b\}$-property.
For edge weighting, affine transformations are more restrictive. An affine map $f(x)=\alpha x+\beta$ preserves the ratio $\frac{f(b_{0})-f(a_{0})}{f(c_{0})-f(b_{0})}=\frac{b_{0}-a_{0}}{c_{0}-b_{0}}$. Thus $\{f(a_{0}),f(b_{0}),f(c_{0})\}$ must have the same spacing ratio as $\{a_{0},b_{0},c_{0}\}$. In particular, $\{1,2,3\}$ can only be mapped to other arithmetic progressions. This limitation means that a single fixed 3-element set cannot generate all 3-element sets by affine transformation.
3. Handling arbitrary 3-element sets directly. Keusch’s theorem guarantees a proper $\{1,2,3\}$-edge weighting for every nice graph. For regular graphs, we extend this result to arbitrary 3-element set $\{a,b,c\}$.
For arithmetic progressions $\{a,b,c\}$ (ba=cbb-a=c-b), scaling reduces the problem to $\{1,2,3\}$, which is covered by Keusch’s theorem.
For non-arithmetic 3-element set, we first apply an affine transformation to obtain an equivalent 3-element set of the form $\{-d_{1},0,d_{2}\}$ with $0<d_{1}<d_{2}$. Theorem 1 then constructs a proper $\{-d_{1},0,d_{2}\}$-edge weighting. The construction works uniformly for all $d_{1},d_{2}$ satisfying $0<d_{1}<d_{2}$, thereby covering all non-arithmetic 3-element sets.
4. Two-phase construction. Theorem 1 is proved via a two-phase procedure.
Phase 1 (Global separation). Using matchings between $I_{0}$ and higher layers, and among the higher layers themselves, we establish the weighted degree separation $d_{w}(x)<d_{w}(z)$ for all $v\in I_{0}$ and $z\in\cup_{i\geq 2}I_{i}$, and $d_{w}(v)<d_{w}(u)$ for all $v\in I_{i}$ and $u\in I_{j}$ with $1\leq i<j$. Remaining conflicts occur only in the bipartite subgraph $G[I_{0},I_{1}]$.
Phase 2 (Local conflict resolution). We classify vertices in $I_{1}$ that are incident to conflicts into three types based on their number of neighbors in $I_{0}$: $X_{3^{+}}$ (at least three neighbors), $X_{2}$ (exactly two neighbors), and $X_{1}$ (exactly one neighbor). Each type is processed with a specific weighting scheme that eliminates conflicts while preserving the separation achieved in Phase 1.
5. Consequences. The two-phase construction shows that regular graphs admit proper 3-edge-weightings for all sets of distinct real numbers, not only for arithmetic progressions. This extends the conclusion of the 1-2-3 Conjecture from the fixed set $\{1,2,3\}$ to arbitrary weight sets of size three, for the class of regular graphs.
Organization
Section 2 is devoted to the proof of Theorem 1. The proof is divided into two main phases: the initial weighting phase (Phase 1) and the conflict-resolution phase (Phase 2).
2. Proof of Theorem 1
We will use the following two lemmas.
Lemma 1.
[2] Each $d$-degenerate graph is $(d+1)$-colorable.
Lemma 2.
[4] Let $G=(V,E)$ be a graph, $I$ be a maximum independent set of $G$, and $I^{\prime}$ be an independent set of $G[V\setminus I]$. Then the bipartite graph $G[I,I^{\prime}]$ admits a matching that saturates $I^{\prime}$.
Lemma 3.
Let $G=(I_{0}\cup I_{1},E)$ be a bipartite graph with minimum degree at least three, and $d_{1},d_{2}$ be two real numbers with $0<d_{1}<d_{2}$. Then $G$ admits a proper $\{-d_{1},0,d_{2}\}$-edge weighting.
Proof.
We would define an edge weighting such that each vertex in $I_{1}$ has weighted degree in $\{-3d_{1},d_{2}\}$ and each one in $I_{0}$ has weighted degree distinct from that in $\{-3d_{1},d_{2}\}$.
In the beginning, every edge is weighted with 0. Let $I_{1}=\{x_{i}:i=1,2,\ldots,p\}$.
First, label three of the edges incident to $x_{1}$ with $d_{1}$. Then $d_{w}(x_{1})=-3d_{1}$, and for each $y\in I_{0}$, $d_{w}(y)\in\{0,-d_{1}\}$; note that $\{0,-d_{1}\}$ is disjoint from $\{-3d_{1},d_{2}\}$.
Second, for $i=2,3,\ldots,p$, suppose:
(i) for each $j\geq i$, no edge incident to $x_{j}$ is weighted;
(ii) for each vertex $y$ in $I_{0}$, $d_{w}(y)\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1},d_{2}\}$;
(iii) for each $j<i$, $d_{w}(x_{j})\in\{-3d_{1},d_{2}\}$.
Case 1: If $x_{i}$ has a neighbor $y^{\ast}$ with $d_{w}(y^{\ast})\in(-\infty,-d_{1})\setminus\{-d_{2}-3d_{1}\}$, let $w(x_{i}y^{\ast})=d_{2}$. After this assignment, $d_{w}(x_{i})=d_{2}$, and the weighted degree of $y^{\ast}$ increases by $d_{2}$. Since before the assignment $d_{w}(y^{\ast})<-d_{1}$, after adding $d_{2}$ we have $d_{w}(y^{\ast})<d_{2}-d_{1}$. Moreover, because originally $d_{w}(y^{\ast})\neq-d_{2}-3d_{1}$, the new value satisfies $d_{w}(y^{\ast})\neq-3d_{1}$. Thus $d_{w}(y^{\ast})\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$, and one can verify that conditions (i), (ii), (iii) are maintained for $i+1$.
Case 2: If each neighbor of $x_{i}$ has a weighted degree in $d_{1},d_{2}-d_{1})\cup\{-d_{2}-3d_{1}\}$, let $y_{1},y_{2},y_{3}$ be three of them. Let $w(x_{i}y_{1})=w(x_{i}y_{2})=w(x_{i}y_{3})=-d_{1}$. Then $d_{w}(x_{i})=-3d_{1}$ and each neighbor of $x_{i}$ has a weighted degree in $2d_{1},d_{2}-d_{1})\cup\{-d_{2}-3d_{1},-d_{2}-4d_{1}\}$ (this set is disjoint with $\{-3d_{1},d_{2}\}$), and one can check that (i), (ii), (iii) hold for $i+1$.
Then finally, each vertex in $I_{0}$ has weighted degree in $(-\infty,d_{2}-d_{1})\setminus\{-3d_{1},d_{2}\}$, while each one in $I_{1}$ has weighted degree in $\{-3d_{1},d_{2}\}$, and we obtain a proper $\{-d_{1},0,d_{2}\}$-edge weighting. ∎
Proof of Theorem 1.
Let $G=(V,E)$ be a $k$-regular graph with $k\geq 3$.
Partition. We partition $V$ into independent sets as follows. Choose a maximum independent set $I_{0}$ in $G$. For $i\geq 1$, let $I_{i}$ be a maximum independent set in $G_{i}$, where $G_{i}=G[V\setminus(\cup_{j<i}I_{j})]$. Let $\ell$ be the maximum integer such that $I_{\ell}\neq\emptyset$. Then $I_{\ell}$ is an independent set, and $1\leq\ell\leq k+1$ since $G$ is nonempty and each vertex in $I_{\ell}$ has a neighbor in $I_{i}$ for each $i<\ell$. If $\ell=1$, the conclusion holds from Lemma 3. So suppose $2\leq\ell\leq k+1$.
We will construct a proper $\{-d_{1},0,d_{2}\}$-edge weighting $w$ where construction proceeds in two phases. In the weighting process, for a vertex $v$ in $I_{i}$ ($i\in[1,\ell]$), we say $v$ is:
  • Type I if $d_{w}(v)=(i-1)d_{2}$, and $v$ is incident to exactly (i1)(i-1) edges of weight $d_{2}$ connecting $v$ to $\cup_{j\geq 1}I_{j}$;
  • Type II if $d_{w}(v)=id_{2}-d_{1}$, and $v$ is incident to exactly $i$ edges of weight $d_{2}$ connecting $v$ to $\cup_{j\geq 1}I_{j}$ and one edge of weight $d_{1}$ connecting $v$ to $I_{0}$;
  • Type III if $v\in I_{i}$ for some $i\geq 2$, $d_{w}(v)=id_{2}-2d_{1}$, and $v$ is incident to exactly $i$ edges of weight $d_{2}$ connecting $v$ to $\cup_{j\geq 1}I_{j}$ and two edges of weight $d_{1}$: one connecting $v$ to $I_{0}$ and the other connecting $v$ to a vertex in $I_{1}$;
  • hungry if $v\in I_{i}$ for some $i\geq 2$, $d_{w}(v)=jd_{2}$ for some $j\leq i-2$ and $v$ is incident to exactly $j$ edges of weight $d_{2}$ connecting $v$ to $\cup_{j\geq 1}I_{j}$;
Goal. We will make sure $w$ satisfies the following conditions:
  • (1)
    for each vertex $v$ in $I_{i}$ with $i\geq 2$, $d_{w}(v)\in\{(i-1)d_{2},id_{2}-2d_{1},id_{2}-d_{1}\}$ (that is, $v$ is Type I, II or III);
  • (2)
    for each $x$ in $I_{1}$, $d_{w}(x)\in\{-3d_{1},-d_{1},0,d_{2}-d_{1}\}$;
  • (3)
    for each $v$ in $I_{0}$, $d_{w}(v)<d_{2}-d_{1}$;
  • (4)
    for each vertex $x$ in $I_{1}$ with $d_{w}(x)=-3d_{1}$, $x$ has at least two neighbors in $I_{0}$, and every neighbor $y$ of $x$ in $I_{0}$ satisfies $d_{w}(y)\neq-3d_{1}$;
  • (5)
    for each vertex $x$ in $I_{1}$ with $d_{w}(x)=-d_{1}$, $x$ has a unique neighbor $y_{x}$ in $I_{0}$, and $d_{w}(y_{x})=0$;
  • (6)
    for each vertex $x$ in $I_{1}$ with $d_{w}(x)=0$, every neighbor $y$ of $x$ in $I_{0}$ satisfies $d_{w}(y)\neq 0$;
Note that in (1), $\max\{(i-1)d_{2},id_{2}-2d_{1}\}<id_{2}-d_{1}$ because $d_{1}<d_{2}$. Moreover, for $1\leq i<j$, one has $id_{2}-d_{1}<\min\{(j-1)d_{2},jd_{2}-2d_{1}\}$. So (1), (2), (3) ensure there is no conflict in $G[\cup_{i\geq 1}I_{i}]$, and here is no conflict between $I_{0}$ and $\cup_{i\geq 2}I_{i}$. Clearly, (3), (4), (5), (6) ensure there is no conflict between $I_{0}$ and $I_{1}$.
The construction proceeds in two phases. Phase 1 ensures condition (1), (2),(3) hold; Phase 2 ensures conditions (4), (5), (6) hold, with (1), (2),(3) remaining valid.
In the beginning, every edge is weighted with 0.
Phase 1: Initial weighting
In this phase, we will make sure:
  • (1’)
    every vertex in $\cup_{i\geq 1}I_{i}$ is Type I or II;
  • (2’)
    for each $v$ in $I_{0}$, $d_{w}(v)\leq 0$;
  • (3’)
    no Type I vertex in $I_{1}$ has a Type I neighbor in $\cup_{i\geq 2}I_{i}$.
Step 1 By Lemma 2, there exists a matching $M_{\ell-1,\ell}$ in the bipartite graph $G[I_{\ell-1},I_{\ell}]$ that saturates $I_{l}$. Label each edge in $M_{\ell-1,\ell}$ with $d_{2}$. Then when $\ell=2$, each vertex in $I_{2}$ is Type I (has weighted degree $d_{2}=(2-1)d_{2}$); when $\ell\geq 3$, each vertex in $I_{\ell}$ is hungry (has weighted degree $d_{2}\leq(\ell-2)d_{2}$). Let $I^{\ast}_{\ell-1}$ be the set of vertices in $I_{\ell-1}$ that are saturated by $M_{\ell-1,\ell}$. Then the following holds.
  • When $\ell=3$, each vertex in $I^{\ast}_{2}$ is Type I (has weighted degree $d_{2}=(2-1)d_{2}$), each vertex in $I_{2}\setminus I^{\ast}_{2}$ is hungry (has weighted degree $0\leq(2-2)d_{2}$), and each (hungry) vertex $u$ in $I_{3}$ is incident to an edge of weight $d_{2}$ whose other endpoint lies in $I_{2}$.
  • When $\ell\geq 4$, each vertex in $I_{\ell-1}$ is hungry (has weighted degree at most $d_{2}\leq[(\ell-1)-2]d_{2}$), and each (hungry) vertex $u$ in $I_{\ell}$ is incident to an edge of weight $d_{2}$ whose other endpoint lies in $I_{\ell-1}$.
  • When $\ell=2$, by Lemma 2, there exists a matching $M^{\ast}_{0,\ell-1}$ in the bipartite graph $G[I_{0},I^{\ast}_{1}]$ that saturates $I^{\ast}_{1}$; label each edge in $M^{\ast}_{0,1}$ with $d_{1}$, which makes each vertex in $I^{\ast}_{1}$ Type II (has weighted degree $d_{2}-d_{1}$), and each one in $I_{1}\setminus I^{\ast}_{1}$ Type I (has weighted degree $0=(1-1)d_{2}$).
Then when $\ell=2$, one can check that (1’), (2’) already hold.
Step 2 When $\ell\geq 3$, we process the indices $i=\ell-2,\ell-3,\ldots,1$ in decreasing order. At the start of step $i$, we assume:
  • (A)
    no edge incident to $\cup_{1\leq j\leq i}I_{j}$ has been weighted;
  • (B)
    every vertex in $I_{\alpha}$ for each $\alpha\geq i+1$ is Type I, Type II or hungry;
  • (C)
    for $\alpha\geq i+2$, if $I_{\alpha}$ has hungry vertices, then: for every hungry vertex $u$ in $I_{\alpha}$ and every $\beta\in[i+1,\alpha-1]$, $u$ is incident to at least one edge of weight $d_{2}$ whose other endpoint lies in $I_{\beta}$.
(The starting point of this inductive hypothesis holds true after Step 1.)
Let $E_{i+1}$ be the set of edges with weight 0 in the subgraph $G[\cup_{j\geq i+1}I_{j}]$. If $E_{i+1}\neq\emptyset$, we process the edges in $E_{i+1}$ one by one. When considering an edge $e$, if both endpoints are hungry, we set $w(e)=d_{2}$; otherwise we leave $e$ unweighted. After processing all edges in $E_{i+1}$, let $U_{i+1}$ be the set of hungry vertices in $\cup_{j\geq i+1}I_{j}$. By construction, if an edge has both endpoints in $U_{i+1}$, then it has weight $d_{2}$; hence the induced subgraph $G[U_{i+1}]$ (if nonempty) has all edges weighted with $d_{2}$.
Take an arbitrary vertex $u\in U_{i+1}$, and suppose $u\in I_{\alpha}$ for some $\alpha\in[i+1,\ell]$. Since $u$ is hungry, $d_{w}(u)=jd_{2}$ for some $j\leq\alpha-2$. By assumption (C), for each $\beta\in[i+1,\alpha-1]$, $u$ is incident to at least one edge of weight $d_{2}$ with other endpoint in $I_{\beta}$. There are $(\alpha-1)-(i+1)+1=\alpha-i-1$ such indices $\beta$. Thus at most $(\alpha-2)-(\alpha-i-1)=i-1$ edges of weight $d_{2}$ incident to $u$ can have their other endpoint in $\cup_{j\geq\alpha+1}I_{j}$. Recall that $G[U_{i+1}]$ has all edges weighted with $d_{2}$. So within the subgraph $G[U_{i+1}]$, the vertex $u$ has at most i1i-1 neighbors in $U_{i+1}\cap(\cup_{j\geq\alpha+1}I_{j})$. Since $I_{\alpha}$ is an independent set and $u$ is arbitrary, it follows that $G[U_{i+1}]$ is (i1)(i-1)-degenerate. By Lemma 1, $U_{i+1}$ can be partitioned into at most $i$ independent sets in $G[U_{i+1}]$. Each such independent set is also independent in $G$. Hence we can write $U_{i+1}=J_{1}\cup\cdots\cup J_{r}$ with $r\leq i$, where each $J_{t}$ is independent in $G$.
By Lemma 2, for each $t\in[1,r]$ there exists a matching $M^{\prime}_{i,t}$ in the bipartite graph $G[I_{i},J_{t}]$ that saturates $J_{t}$. For every edge $e$ in $\bigcup_{t=1}^{r}M^{\prime}_{i,t}$, set $w(e)=d_{2}$. At this point, each vertex $v\in I_{i}$ is incident to at most $i$ edges of weight $d_{2}$ from these matchings, so $d_{w}(v)=jd_{2}$ for some $j\leq r\leq i$.
Let $\overline{I}_{i}\subseteq I_{i}$ be the set of vertices with weighted degree exactly $id_{2}$. Applying Lemma 2 again, there exists a matching $\overline{M}_{0,i}$ in $G[I_{0},\overline{I}_{i}]$ that saturates $\overline{I}_{i}$. For each edge $e\in\overline{M}_{0,i}$, set $w(e)=-d_{1}$. Then each vertex in $\overline{I}_{i}$ becomes Type II, and each one in $I_{i}\setminus\overline{I}_{i}$ is either Type I or hungry.
After these assignments, the following hold:
  • (A’)
    no edge incident to $\cup_{1\leq j\leq i-1}I_{j}$ has been weighted;
  • (B’)
    every vertex in $I_{\alpha}$ for each $\alpha\geq i$ is Type I, Type II or hungry;
  • (C’)
    for $\alpha\geq i+1$, if $I_{\alpha}$ has hungry vertices, then: for every hungry vertex $u$ in $I_{\alpha}$ and every $\beta\in[i,\alpha-1]$, $u$ is incident to at least one edge of weight $d_{2}$ whose other endpoint lies in $I_{\beta}$.
Indeed, (B’) holds because: if $v\in\cup_{j\geq i+1}I_{j}$ was already Type I or II after step $i+1$, it remains Type I or II; if $u\in U_{i+1}$ received an edge of weight $d_{2}$ from some $M^{\prime}_{i,t}$, it now becomes Type I or remains hungry (the extra $d_{1}$ edges to $I_{0}$ do not affect the condition for vertices in $U_{i+1}$); if $v\in I_{i}$, then either it received at most i1i-1 edges of weight $d_{2}$ from matchings $M^{\prime}_{i,t}$’s (in which case it becomes Type I or remains hungry), or received exactly $i$ edges of weight $d_{2}$ and exactly one edge of weight $d_{1}$ from $\overline{M}_{0,i}$ (in which case it becomes Type II).
Proceeding in this way for $i=\ell-2,\ell-3,\ldots,1$, we obtain at the end of Phase 1 that every vertex in $\cup_{\alpha\geq 1}I_{\alpha}$ is Type I or II. Since at the end, on one hand, each $v$ in $I_{1}$ has weighted degree $d_{2}-d_{1}$ (Type II) or 0 (Type I). On the other hand, if some $u$ in $I_{\alpha}$ ($\alpha\geq 2$) is hungry, it should satisfies (C’): $u$ is incident to at least one edge of weight $d_{2}$ with other endpoint in $I_{\beta}$ for each $\beta\in[1,\alpha-1]$ (at the end $i=1$). That is, $d_{w}(u)\geq(\alpha-1)d_{2}$ and $u$ is not hungry, a contradiction. Consequently, condition (1’) holds and there is no conflict within the subgraph $G[\cup_{i\geq 1}I_{i}]$. Moreover, every vertex $v\in I_{0}$ satisfies $d_{w}(v)\leq 0$ at this stage and (2’) holds.
For (3’), if there exists a Type I vertex $v$ in $I_{1}$ that has a Type I neighbor $u$ in some $I_{i}$ ($i\geq 2$), let $e_{v}$ and $e_{u}$ be the edges joining $v$ and $u$, respectively, to $I_{0}$ (recall that each vertex in $\cup_{j\geq 1}I_{j}$ has a neighbor in $I_{0}$ since $I_{0}$ is a maximum independent set). Clearly, $e_{v},e_{u},uv$ have not been weighted since $v,u$ are Type I (moreover, $v$ is Type I and in $I_{1}$). Set $w(e_{v})=w(e_{u})=-d_{1}$ and set $w(uv)=d_{2}$. Then both $u,v$ become Type II (so condition (1’) remains valid), and every vertex in $I_{0}$ still has weighted degree at most 0 (so condition (2’) remains valid). Repeat this step until no Type I vertex in $I_{1}$ has a Type I neighbor in $\cup_{i\geq 2}I_{i}$, and (3’) holds.
Phase 2: Conflict resolution
For $i=0,1$, let $I_{i,0}=\{v\in I_{i}:d_{w}(v)=0\}.$ After Phase 1, by (1’) and (2’), the only possible conflicts in $G$ are those belonging to the bipartite subgraph $G[I_{0,0},I_{1,0}]$. If $G[I_{0,0},I_{1,0}]$ is empty, we are done. Assume therefore that it is nonempty. For each vertex $x\in I_{1,0}$, choose a neighbor $y_{x}$ (if exists) of $x$ in $I_{0,0}$.
Remark on dynamics: During Phase 2, the sets $I_{1,0}$, $I_{0,0}$, and the classifications based on them will change as we assign new weights. We will carefully track these changes to ensure the algorithm terminates.
At this point, every vertex in $I_{1,0}$ has all its neighbors either in $I_{0}$ or in Type II vertices of $\cup_{i\geq 2}I_{i}$ by (1’) and (3’).
We classify the vertices of $I_{1,0}$ according to their number of neighbors in $I_{0}$:
$I_{1,0}=X_{2^{+}}\cup X_{1}=X_{3^{+}}\cup X_{2}\cup X_{1},$
where $X_{i^{+}}=\{x\in I_{1,0}:|N(x)\cap I_{0}|\geq i\}$ and $X_{i}=\{x\in I_{1,0}:|N(x)\cap I_{0}|=i\}$ for $i\geq 1$.
Now consider the bipartite subgraph $H^{\prime}:=G[X_{2}\cup X_{1},\cup_{i\geq 2}I_{i}]$. On one hand, for any $x\in X_{2}\cup X_{1}$:
  • If $x\in X_{2}$, then $|N(x)\cap I_{0}|=2$, so $x$ has at least k2k-2 neighbors in $\cup_{i\geq 2}I_{i}$ (since all neighbors are either in $I_{0}$ or Type II vertices of $\cup_{i\geq 2}I_{i}$).
  • If $x\in X_{1}$, then $|N(x)\cap I_{0}|=1$, so $x$ has at least $k-1\geq k-2$ neighbors in $\cup_{i\geq 2}I_{i}$.
For any Type II vertex $z\in\cup_{i\geq 2}I_{i}$, by definition it is incident to exactly one edge of weight $d_{1}$ (to a vertex in $I_{0}$) and exactly $i$ edges of weight $d_{2}$ for some $i\geq 2$. Since all edges incident to vertices in $I_{1,0}$ currently have weight 0, any edge of weight $d_{2}$ incident to $z$ must connect $z$ to a vertex not in $I_{1,0}$. Thus $z$ has at least one neighbor not in $I_{1,0}$ via an edge of weight $d_{2}$, plus the edge of weight $d_{1}$ to $I_{0}$. Consequently, $z$ can be adjacent to at most k2k-2 vertices in $I_{1,0}$ (so to at most k2k-2 vertices in $X_{2}\cup X_{1}$).
Thus, in $H^{\prime}$, every vertex in $X_{2}\cup X_{1}$ has degree at least k2k-2, and every vertex in $\cup_{i\geq 2}I_{i}$ has degree at most k2k-2. To verify Hall’s condition for a matching saturating $X_{2}\cup X_{1}$, take any subset $S\subseteq X_{2}\cup X_{1}$. The number of edges incident to $S$ in $H^{\prime}$ is at least |S|(k2)|S|(k-2). Since each vertex in the neighborhood $N_{H^{\prime}}(S)$ has degree at most k2k-2, we have
$|S|(k-2)\leq e(S,N_{H^{\prime}}(S))\leq|N_{H^{\prime}}(S)|(k-2).$
Dividing by k2>0k-2>0 (as $k\geq 3$) gives $|S|\leq|N_{H^{\prime}}(S)|$. Hence Hall’s condition holds, and there exists a matching $M$ in $H^{\prime}$ that saturates $X_{2}\cup X_{1}$. Write $M=M_{1}\cup M_{2}$, where $M_{i}$ saturates $X_{i}$ for $i=1,2$.
We will resolve the conflicts involving $X_{2^{+}}$ and $X_{1}$ separately.
1. Resolving conflicts incident to $X_{2^{+}}$.
Let $X^{\ast}_{2^{+}}\subseteq X_{2^{+}}$ be the set of vertices that have a neighbor in $I_{0,0}$. If $X^{\ast}_{2^{+}}=\emptyset$, then we are done.
We process vertices from $X_{2^{+}}$ iteratively. Each iteration removes one vertex from $X_{2^{+}}$ (and hence from $X^{\ast}_{2^{+}}$). Although $X^{\ast}_{2^{+}}$ may expand due to vertices re-entering $I_{0,0}$ in Case 1, the set $X_{2^{+}}$ itself strictly decreases in size with each iteration. Therefore, the loop terminates after at most $|X_{2^{+}}|$ iterations.
Let $i=1$. While $X^{\ast}_{2^{+}}\neq\emptyset$, we select a vertex $x_{i}$ from $X^{\ast}_{2^{+}}$ (arbitrarily for $i=1$, and from the updated $X^{\ast}_{2^{+}}$ for $i>1$) and handle it according to the cases below. After processing $x_{i}$, it is removed from $I_{1,0}$ (since its weighted degree becomes non-zero), and consequently from $X_{2^{+}}$ and $X^{\ast}_{2^{+}}$ (this process will gradually eliminates vertices from $X_{2^{+}}$ until $X^{\ast}_{2^{+}}$ becomes empty). We then update $I_{1,0}$, $X_{2^{+}}$, $X^{\ast}_{2^{+}}$ accordingly.
Before processing $x_{i}$, we assume:
  • (i)
    no edge incident to unprocessed vertices in $X_{2^{+}}$ has been weighted;
  • (ii)
    for every $y\in I_{0}$, $d_{w}(y)<d_{2}-d_{1}$;
  • (iii)
    for each previously processed vertex $x_{j}$ ($j<i$), one of the following items holds:
    • (i’)
      $d_{w}(x_{j})=d_{2}-d_{1}$;
    • (ii’)
      $d_{w}(x_{j})=-3d_{1}$ and every neighbor $y$ of $x_{j}$ in $I_{0}$ satisfies $d_{w}(y)\neq-3d_{1}$.
(For $i=1$, conditions (i) and (ii) hold; no vertex has been processed and so condition (iii) holds vacuously.)
We distinguish three cases.
Case 1: $x_{i}\in X_{2^{+}}$, and $x_{i}$ has a neighbor $y^{\ast}\in I_{0}$ with $d_{w}(y^{\ast})\in(-\infty,-d_{1})\setminus\{-d_{2}-3d_{1}\}$. Set $w(x_{i}y^{\ast})=d_{2}$ and $w(x_{i}y_{x_{i}})=-d_{1}$. After these assignments, $d_{w}(x_{i})=d_{2}-d_{1}$ (hence condition (iii)(i’) holds for $x_{i}$), $d_{w}(y_{x_{i}})=-d_{1}$, and the weighted degree of $y^{\ast}$ increases by $d_{2}$. Since before the assignment $d_{w}(y^{\ast})<-d_{1}$, after adding $d_{2}$ we have $d_{w}(y^{\ast})<d_{2}-d_{1}$, which fulfills condition (ii). Furthermore, because originally $d_{w}(y^{\ast})\neq-d_{2}-3d_{1}$, we have $d_{w}(y^{\ast})\neq-3d_{1}$. Thus $d_{w}(y^{\ast})\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$, and conditions (ii) and (iii) continue to hold for $i+1$.
Case 2: $x_{i}\in X_{3^{+}}$ and every neighbor of $x_{i}$ in $I_{0}$ has weighted degree in $d_{1},d_{2}-d_{1})\cup\{-d_{2}-3d_{1}\}$. Since $x_{i}\in X_{3^{+}}$, it has at least three neighbors in $I_{0}$. Arbitrarily choose three edges $x_{i}y_{1},x_{i}y_{2},x_{i}y_{3}$ in $G[I_{0},X_{3^{+}}]$ and set $w(x_{i}y_{1})=w(x_{i}y_{2})=w(x_{i}y_{3})=-d_{1}$. This gives $d_{w}(x_{i})=-3d_{1}$, while $d_{w}(y_{1}),d_{w}(y_{2}),d_{w}(y_{3})\in[-2d_{1},d_{2}-2d_{1})\cup\{-d_{2}-4d_{1}\}$. Then every neighbor of $x_{i}$ in $I_{0}$ has weighted degree in $2d_{1},d_{2}-d_{1})\cup\{-d_{2}-4d_{1},-d_{2}-3d_{1}\}$; this set is consistent with (ii), and does not include $3d_{1}$ and so (iii) also holds for $i+1$.
Case 3: $x_{i}\in X_{2}$ and every neighbor of $x_{i}$ in $I_{0}$ has weighted degree in $d_{1},d_{2}-d_{1})\cup\{-d_{2}-3d_{1}\}$. Let $y^{\ast}$ be the unique neighbor of $x_{i}$ in $I_{0}$ other than $y_{x_{i}}$. Let $x_{i}z$ be the edge in $M_{2}$ that saturates $x_{i}$, where $z\in I_{q}$ for some $q\geq 2$. By (1’) and (3’), $z$ is Type II and so $d_{w}(z)=qd_{2}-d_{1}$. Now set $w(x_{i}y_{x_{i}})=w(x_{i}y^{\ast})=w(x_{i}z)=-d_{1}$. This gives $d_{w}(x_{i})=-3d_{1}$, while $d_{w}(y_{x_{i}})=-d_{1}$, $d_{w}(z)=qd_{2}-2d_{1}$ (that is, $z$ becomes Type III and so condition (1) remains valid), and $d_{w}(y^{\ast})\in[-2d_{1},d_{2}-2d_{1})\cup\{-d_{2}-4d_{1}\}$. Then $\{d_{w}(y^{\ast}),d_{w}(y_{x_{i}})\}\subseteq[-2d_{1},d_{2}-d_{1})\cup\{-d_{2}-4d_{1}\}$; this set is consistent with (ii), and does not include $3d_{1}$ and so (iii) also holds for $i+1$.
Remark on the dynamics of $I_{0,0}$: In Case 1, it is possible that $d_{w}(y^{\ast})$ becomes 0 after adding $d_{2}$ (if originally $d_{w}(y^{\ast})=-d_{2}$), placing $y^{\ast}$ back into $I_{0,0}$. However, by assumption (i), all edges incident to $\{x_{j}:j>i\}$ are still unweighted, so any new conflict involving $y^{\ast}$ and $x_{j}$ ($j>i$) will be resolved when $x_{j}$ is processed. Moreover, $y^{\ast}$ cannot conflict with already processed vertices $x_{p}$ ($p<i$) because those have weighted degrees either $d_{2}-d_{1}$ or $3d_{1}$, both different from 0. For cases 2 and 3, all modified weighted degrees in $I_{0}$ become negative; so $I_{0}$-vertices cannot re-enter $I_{0,0}$.
Remark on the dynamics of $X^{\ast}_{2^{+}}$: Since (in Case 1) $I_{0,0}$ may expand, $X^{\ast}_{2^{+}}$ may also expand. However, $X^{\ast}_{2^{+}}$ is included in $X_{2^{+}}$, and $X_{2^{+}}$ decreases at each step. So the recursive process will terminate with $X_{2^{+}}=\emptyset$.
After handling $x_{i}$, update $I_{0,0},X_{2^{+}}$ and $X^{\ast}_{2^{+}}$ respectively. Note that $x_{i}$ and all previously processed vertices $x_{p}$ ($p<i$) are no longer in $X^{\ast}_{2^{+}}$ (no longer in $X_{2^{+}}$ either), since they either have weighted degree $d_{2}-d_{1}$ (which differs from 0), or have weighted degree $3d_{1}$ while all their neighbors in $I_{0}$ have weighted degrees different from $3d_{1}$. If $X^{\ast}_{2^{+}}=\emptyset$, we are done. If $X^{\ast}_{2^{+}}\neq\emptyset$ let $x_{i+1}\in X^{\ast}_{2^{+}}$ and the assumptions (i)–(iii) hold for $i+1$. Proceeding and update $X^{\ast}_{2^{+}}$ repeatedly, we eventually eliminate all conflicts in $G[I_{0},X_{2^{+}}]$. Moreover, (1), (2), (3) remain valid currently. Hence no conflict remains incident to $X_{2^{+}}$.
2. Resolving conflicts incident to $X_{1}$.
After resolving conflicts for $X_{2^{+}}$, we update $I_{0,0}$. Note that $X_{1}$ remains unchanged. Let $X^{\ast}_{1}\subseteq X_{1}$ be the set of vertices in the current $X_{1}$ that have a neighbor in (updated) $I_{0,0}$. If $X^{\ast}_{1}=\emptyset$, we are finished.
Assume $X^{\ast}_{1}\neq\emptyset$ and let $M^{\ast}_{1}$ be the subset of $M_{1}$ consisting of edges incident to vertices in $X^{\ast}_{1}$. Recall that each edge in $M^{\ast}_{1}$ connects a vertex $x\in X^{\ast}_{1}$ to a Type II vertex $z\in\cup_{i\geq 2}I_{i}$, and each edge in $M^{\ast}_{1}$ is unweighted. Assign weight $d_{1}$ to every edge in $M^{\ast}_{1}$. After the assignment:
  • $d_{w}(x)$ changes from 0 to $d_{1}$;
  • $d_{w}(y_{x})$ remains 0 (since edge $xy_{x}$ is not in $M^{\ast}_{1}$ and keeps weight 0);
  • $d_{w}(z)$ decreases by $d_{1}$ for the corresponding $z\in\cup_{i\geq 2}I_{i}$ which becomes Type III.
This implies that condition (5) holds, and (1), (2), (3) remain valid. Note that, at present, both $X^{\ast}_{2^{+}}$ and $X^{\ast}_{1}$ are empty, and so (6) holds. Condition (4) holds because: vertices with $d_{w}(x)=-3d_{1}$ come from Case 2 or Case 3. In Case 2, $x\in X_{3^{+}}$ has at least three neighbors in $I_{0}$; in Case 3, $x\in X_{2}$ has exactly two neighbors in $I_{0}$. In both cases, all neighbors $y$ of $x$ in $I_{0}$ satisfy $d_{w}(y)\neq-3d_{1}$ by construction. Thus, $w$ is a proper $\{-d_{1},0,d_{2}\}$-edge weighting of $G$. This completes the proof of Theorem 1. ∎
3. Concluding Remarks
Theorem 1 and Corollary 2 establish that every regular graph is universally 3-edge-weightable. The independent-set decomposition technique, originally developed for total weightings, proves adaptable to edge-weightings with arbitrary 3-element set of weights.
Our work suggests that regularity provides a structural framework within which arbitrary three-element weight set can be accommodated. The conflict-resolution scheme developed here may find applications in other weighting and labeling problems. The two-phase approach—global matching adjustments followed by local conflict resolution—offers a template for handling parameter-dependent constructions in combinatorial graph theory.
Future work. Several natural questions arise from our result:
  • Does universal 3-edge-weightability extend to all nice graphs? (Problem 1.1)
  • What other graph classes (beyond regular graphs) admit this property?
  • Can the techniques be extended to study edge-weight choosability problems?
Acknowledgments
The first author is supported by NSFC (No. 11701195) and Fundamental Research Funds for the Central Universities (No. ZQN-904).

### v2

@@
-Universal 3-Edge-Weightability of Regular Graphs
+Regular graphs are universally 3-edge-weightable
 Abstract.
-A graph $G=(V,E)$ is universally $3$-edge-weightable if for every set $\{a,b,c\}$ of three distinct real numbers, there exists an edge-weighting $w:E\to\{a,b,c\}$ such that $d_{w}(u)\neq d_{w}(v)$ for all $uv\in E$.
-The 1-2-3 Conjecture, recently settled by Keusch, ensures that all nice graphs admit a proper $\{1,2,3\}$-edge-weighting.
-For regular graphs, this implies proper edge-weightings for arithmetic progressions $\{a,b,c\}$ via affine transformation, but not for general triples due to the spacing-ratio preservation property of affine maps.
-We prove that every $k$-regular graph with $k\geq 2$ is universally $3$-edge-weightable.
-This establishes proper edge-weightings for all triples $\{a,b,c\}$, not just arithmetic progressions, showing that regular graphs satisfy a strictly stronger weighting property than what follows from the 1-2-3 Conjecture alone.
-The proof adapts the independent-set decomposition framework to edge-weightings and introduces a two-phase conflict-resolution scheme that operates independently of the specific choice of $\{a,b,c\}$.
+A graph is universally $k$-edge-weightable if for every $k$-element set $Q\subset\mathbb{R}$, it admits a proper $Q$-edge weighting.
+The settled 1-2-3 conjecture implies that for any arithmetic progression $\{a,b,c\}$, every nice regular graph has a proper $\{a,b,c\}$-edge weighting.
+We prove that this remains valid for all 3-element set $\{a,b,c\}$ with $c-b\neq b-a$.
+Consequently, every nice regular graph is universally $3$-edge-weightable.
 Key words and phrases:
@@
 All graphs considered are simple, finite, and undirected.
-A graph is nice if none of its components is isomorphic to $K_{2}$.
-For a graph $G=(V,E)$ and a subset $Q\subset\mathbb{R}$, a mapping $w:E\to Q$ is called a $Q$-edge weighting of $G$; the weighted degree of a vertex $v$ is defined as $d_{w}(v)=\sum_{e\ni v}w(e)$; an edge $e=uv$ is a conflict if $d_{w}(u)=d_{w}(v)$.
-If there is no conflict under $w$, then $w$ is called proper.
-We say $G$ is $Q$-edge-weightable if it admits a proper $Q$-edge weighting, and call $G$ universally $k$-edge-weightable if it is $Q$-edge weightable for every $k$-element subset $Q\subset\mathbb{R}$.
-The 1-2-3 Conjecture, posed by Karoński, Łuczak, and Thomason [5], asserted that every nice graph is $\{1,2,3\}$-edge-weightable.
-After remaining open for two decades, this conjecture was finally confirmed by Keusch [6] in 2024.
-The property of being $\{1,2,3\}$-edge-weightable corresponds precisely to universal 3-edge-weightability restricted to the specific set $\{1,2,3\}$.
-For regular graphs, the 1-2-3 result can be extended via an affine transformation to show that they are $\{a,b,c\}$-edge-weightable for any arithmetic progression $\{a,b,c\}$ (where ba=cbb-a=c-b).
-However, this approach fails for general (non-arithmetic) three-element sets, leaving open the question of whether regular graphs are universally 3-edge-weightable.
-Various strengthenings of the 1-2-3 Conjecture have been investigated.
-A graph is edge-weight $k$-choosable if for any assignment $L$ that gives each edge $e$ a list $L(e)$ of $k$ real numbers, there exists a proper edge weighting $w$ with $w(e)\in L(e)$ for all $e\in E$.
-The list version of the 1-2-3 Conjecture [1] asks whether every nice graph is edge-weight $3$-choosable.
-A stronger concept is $(1,3)$-choosability [7]: a graph is $(1,3)$-choosable if, for an arbitrary vertex weighting $f:V\rightarrow\mathbb{R}$ and any assignment of $3$-element lists to edges, there exists an edge weighting $w$ using weights from the respective lists such that $f(u)+d_{w}(u)\neq f(v)+d_{w}(v)$ for every edge $uv\in E$.
+For notation or terminology not defined here, we refer to [1].
+A graph is nice if it contains no $K_{2}$ component.
+For $G=(V,E)$ and $Q\subset\mathbb{R}$, a $Q$-edge weighting is a mapping $w:E\to Q$.
+When $Q=\{1,2,\ldots,k\}$, it is called a $k$-edge weighting.
+The weighted degree of $v$ is $d_{w}(v)=\sum_{e\ni v}w(e)$.
+An edge $uv$ is a conflict if $d_{w}(u)=d_{w}(v)$; $w$ is proper if it has no conflicts.
+The 1-2-3 conjecture, posed by Karoński, Łuczak, and Thomason [2], asserts that every nice graph has a proper $\{1,2,3\}$-edge weighting.
+This conjecture is a local version of the classical–and still active–irregularity strength problem [3, 4, 6, 7, 5, 8, 9].
+Research on the 1-2-3 conjecture has progressed along several fronts.
+Upper bounds for general graphs were successively lowered: from 30 [10], to 16 [11], 13 [12], 5 [13], and finally to 3 [14] after an intermediate improvement to 4 [15].
+For specific classes, it was verified for dense graphs [16], $d$-regular graphs with large $d$ [17], graphs with $\delta=\Omega(\log\Delta)$ [18], and shown to hold asymptotically almost surely for random graphs [11].
+Complexity results include NP-completeness for deciding proper $\{1,2\}$-edge weightings [19] and a polynomial-time algorithm for bipartite graphs [20].
+The conjecture was affirmatively resolved by Keusch [14], leading to the 1-2-3 theorem.
+For more 1-2-3 conjecture related problems, refer to [21, 22, 23, 24, 25, 26, 27] and the references cited there.
+Various strengthenings of the 1-2-3 conjecture have been studied.
+The list version [28] asks whether every nice graph is edge-weight $3$-choosable, i.e.
+for any assignment of $3$-element lists $L(e)$ to edges, there exists a proper edge weighting $w$ with $w(e)\in L(e)$ for all $e$.
+Wong and Zhu [29] conjectured a stronger property called $(1,3)$-choosability, where each vertex receives a fixed weight (from a $1$-element list) and each edge receives a weight from a $3$-element list; the weighted degree of a vertex is then $d_{w}(v)=w(v)+\sum_{e\ni v}w(e)$.
 These properties form a clear hierarchy:
-$(1,3)\text{-choosable}\ \Rightarrow\ 3\text{-choosable}\ \Rightarrow\ \text{universally 3-edge-weightable}\ \Rightarrow\ \{1,2,3\}\text{-edge-weightable}.$
-Progress on the strongest property, $(1,3)$-choosability, has been steady.
-Cao [3] proved that every nice graph is $(1,17)$-choosable, improved by Zhu [8] to $(1,5)$-choosability.
+$(1,3)\text{-choosable}\Rightarrow\text{edge-weight 3-choosable}\Rightarrow\text{the 1-2-3 theorem}.$
+Progress on the strongest property, $(1,3)$-choosability, has been steady: Cao [30] proved that every nice graph is $(1,17)$-choosable, improved by Zhu [31] to $(1,5)$-choosability.
 Consequently, every nice graph is edge-weight $5$-choosable.
@@
 Partial results show that several graph classes are $(1,3)$-choosable, including complete graphs and complete bipartite graphs [28], generalized Halin graphs [32], graphs with maximum average degree less than $\frac{11}{4}$ [33], dense Eulerian graphs [34], and Cartesian products of paths and cycles [35].
-Between the weakest property (settled $\{1,2,3\}$-edge-weightability) and the stronger open problem of $3$-choosability lies the natural question of universal 3-edge-weightability:
-Problem 1.1 (Universal 3-edge-weightability).
-Is every nice graph universally 3-edge-weightable?
-In other words, can the fixed set $\{1,2,3\}$ be replaced by an arbitrary 3-element set?
-Our result (Corollary 2) shows that regular graphs are universally 3-edge-weightable, while Problem 1 asks whether this holds for all nice graphs.
-A related but different line of research concerns total weightings, where weights are assigned to both vertices and edges.
-For this total weighting problem, the 1-2 Conjecture [10] asks whether every nice graph is $\{1,2\}$-total-weightable.
-Notably, for regular graphs a stronger universal 2-total-weightability has been established [4], which serves as motivation for the present work on universal 3-edge-weightability.
-Our contribution
-In this paper, we demonstrate that the independent set decomposition framework in [4] can be adapted from total weighting to edge weighting, and can furthermore handle universal 3-edge-weighting.
-Our main technical result is the following:
+The 1-2-3 theorem guarantees a proper $\{1,2,3\}$-edge weighting for every nice graph.
+A natural weakening of the unsolved $(1,3)$-choosability conjecture and edge-weight $3$-choosability conjecture, is to ask whether this extends to any three-element set $Q\subset\mathbb{R}$, not just $\{1,2,3\}$.
+We call a graph universally $3$-edge-weightable if it admits a proper $Q$-edge weighting for every such $Q$.
+For regular graphs, affine transformation from the 1-2-3 theorem immediately gives a proper $\{a,b,c\}$-edge weighting for any arithmetic progression $\{a,b,c\}$ (where ba=cbb-a=c-b).
+The non-arithmetic case, however, requires genuinely new constructions.
+This paper provides the first such construction, which implies that every regular graph is universally $3$-edge-weightable.
 Theorem 1.
-For any positive real numbers $d_{1},d_{2}$ with $0<d_{1}<d_{2}$, every $k$-regular graph with $k\geq 3$ is $\{-d_{1},0,d_{2}\}$-edge-weightable.
-A simple scaling and shifting argument yields the following corollary from Theorem 1.
+For any positive real numbers $d_{1},d_{2}$ with $0<d_{1}<d_{2}$, every $k$-regular graph with $k\geq 3$ admits a proper $\{-d_{1},0,d_{2}\}$.
 Corollary 2.
@@
 For $k=2$, a $2$-regular graph consists of disjoint cycles.
-Graphs with maximum average degree less than $\frac{11}{4}$ are known to be $(1,3)$-choosable [9].
-Taking the vertex weighting $f\equiv 0$ and giving every edge the same list $\{a,b,c\}$, $(1,3)$-choosability implies the existence of a proper $\{a,b,c\}$-edge weighting.
+Graphs with maximum average degree less than $\frac{11}{4}$ are known to be $(1,3)$-choosable [33], implying the existence of a proper $\{a,b,c\}$-edge weighting.
 Thus we may assume $k\geq 3$.
 It suffices to show that $G$ admits a proper $\{-(b-a),0,c-b\}$-edge weighting $w^{\prime}$.
-Indeed, if we then define $w=w^{\prime}+b$ (adding $b$ to every edge weight), the regularity of $G$ ensures that $w$ is a proper $\{a,b,c\}$-edge weighting.
+Indeed, if we define $w=w^{\prime}+b$ (that is, we add $b$ to every edge weight), the regularity of $G$ ensures that $w$ is a proper $\{a,b,c\}$-edge weighting.
 Consider three cases.
@@
 ∎
-The result of Corollary 2 places regular graphs strictly between the settled 1-2-3 Conjecture and the open $3$-choosability question in the hierarchy of edge-weighting properties.
-Methodological remarks
-Our proof of Theorem 1 adapts the independent-set decomposition framework from total-weightings to edge-weightings while addressing several distinctive challenges.
-1.
-Absence of vertex-weight adjustments.
-In total weighting, vertex weights provide an independent degree of freedom that can be used to resolve conflicts.
-Edge weighting lacks this flexibility.
-To compensate, after partitioning the vertices into special independent sets $I_{0},I_{1},\ldots,I_{\ell}$, we employ the maximum independent set $I_{0}$ as a global regulator.
-We systematically weight $G[\cup_{i\geq 1}I_{i}]$ with $\{0,d_{2}\}$, and assign weight $d_{1}$ to selected edges through matchings between $I_{0}$ and $\cup_{i\geq 1}I_{i}$.
-This creates the separation $d_{w}(v)<d_{w}(u)$ for all $v\in I_{i}$ and $u\in I_{j}$ with $1\leq i<j$, and separation $d_{w}(x)<d_{w}(z)$ for all $v\in I_{0}$ and $z\in\cup_{i\geq 2}I_{i}$.
-The separation prevents conflicts between $I_{0}$ and $G[\cup_{i\geq 2}I_{i}]$, and conflicts in $G[\cup_{i\geq 1}I_{i}]$.
-2.
-Different behavior under affine transformations.
-For total weighting on $k$-regular graphs, any proper $\{a_{0},b_{0}\}$-total weighting $w$ can be affinely transformed into a proper $\{a,b\}$-total weighting $w^{\prime}$.
-The transformation $w^{\prime}\mapsto\frac{b-a}{b_{0}-a_{0}}(w-a_{0})+a$, applied uniformly to vertex and edge weights, yields
-$d_{w^{\prime}}(v)=\frac{b-a}{b_{0}-a_{0}}d_{w}(v)+(k+1)\left(a-\frac{b-a}{b_{0}-a_{0}}a_{0}\right),$
-which preserves properness because the additive term is constant across vertices. Proving the $\{0,1\}$-case therefore suffices for the universal $\{a,b\}$-property.
-For edge weighting, affine transformations are more restrictive.
-An affine map $f(x)=\alpha x+\beta$ preserves the ratio $\frac{f(b_{0})-f(a_{0})}{f(c_{0})-f(b_{0})}=\frac{b_{0}-a_{0}}{c_{0}-b_{0}}$.
-Thus $\{f(a_{0}),f(b_{0}),f(c_{0})\}$ must have the same spacing ratio as $\{a_{0},b_{0},c_{0}\}$.
-In particular, $\{1,2,3\}$ can only be mapped to other arithmetic progressions.
-This limitation means that a single fixed 3-element set cannot generate all 3-element sets by affine transformation.
-3.
-Handling arbitrary 3-element sets directly.
-Keusch’s theorem guarantees a proper $\{1,2,3\}$-edge weighting for every nice graph.
-For regular graphs, we extend this result to arbitrary 3-element set $\{a,b,c\}$.
-For arithmetic progressions $\{a,b,c\}$ (ba=cbb-a=c-b), scaling reduces the problem to $\{1,2,3\}$, which is covered by Keusch’s theorem.
-For non-arithmetic 3-element set, we first apply an affine transformation to obtain an equivalent 3-element set of the form $\{-d_{1},0,d_{2}\}$ with $0<d_{1}<d_{2}$.
-Theorem 1 then constructs a proper $\{-d_{1},0,d_{2}\}$-edge weighting.
-The construction works uniformly for all $d_{1},d_{2}$ satisfying $0<d_{1}<d_{2}$, thereby covering all non-arithmetic 3-element sets.
-4. Two-phase construction. Theorem 1 is proved via a two-phase procedure.
-Phase 1 (Global separation).
-Using matchings between $I_{0}$ and higher layers, and among the higher layers themselves, we establish the weighted degree separation $d_{w}(x)<d_{w}(z)$ for all $v\in I_{0}$ and $z\in\cup_{i\geq 2}I_{i}$, and $d_{w}(v)<d_{w}(u)$ for all $v\in I_{i}$ and $u\in I_{j}$ with $1\leq i<j$.
-Remaining conflicts occur only in the bipartite subgraph $G[I_{0},I_{1}]$.
-Phase 2 (Local conflict resolution).
-We classify vertices in $I_{1}$ that are incident to conflicts into three types based on their number of neighbors in $I_{0}$: $X_{3^{+}}$ (at least three neighbors), $X_{2}$ (exactly two neighbors), and $X_{1}$ (exactly one neighbor).
-Each type is processed with a specific weighting scheme that eliminates conflicts while preserving the separation achieved in Phase 1.
-5.
-Consequences.
-The two-phase construction shows that regular graphs admit proper 3-edge-weightings for all sets of distinct real numbers, not only for arithmetic progressions.
-This extends the conclusion of the 1-2-3 Conjecture from the fixed set $\{1,2,3\}$ to arbitrary weight sets of size three, for the class of regular graphs.
-Organization
-Section 2 is devoted to the proof of Theorem 1. The proof is divided into two main phases: the initial weighting phase (Phase 1) and the conflict-resolution phase (Phase 2).
+This settles the regular case completely. Whether universal $3$-edge-weightability holds for all nice graphs remains open:
+Problem 1.1 (Universal 3-edge-weightability).
+Is every nice graph universally $3$-edge-weightable?
+Theorem 1 is proved in Section 2.
 2. Proof of Theorem 1
-We will use the following two lemmas.
+We will use the following three lemmas.
 Lemma 1.
-[2] Each $d$-degenerate graph is $(d+1)$-colorable.
+[1] Each $i$-degenerate graph $(i\geq 1)$ is $(i+1)$-colorable.
 Lemma 2.
-[4] Let $G=(V,E)$ be a graph, $I$ be a maximum independent set of $G$, and $I^{\prime}$ be an independent set of $G[V\setminus I]$.
-Then the bipartite graph $G[I,I^{\prime}]$ admits a matching that saturates $I^{\prime}$.
+[36] Let $G=(V,E)$ be a graph, $I_{0}$ be a maximum independent set of $G$, and $I_{1}$ be an independent set of $G[V\setminus I_{0}]$.
+Then the bipartite graph $G[I_{0},I_{1}]$ admits a matching saturating $I_{1}$.
 Lemma 3.
-Let $G=(I_{0}\cup I_{1},E)$ be a bipartite graph with minimum degree at least three, and $d_{1},d_{2}$ be two real numbers with $0<d_{1}<d_{2}$.
+Let $G=(I_{0}\cup I_{1},E)$ be a bipartite graph with $d(v)\geq 3$ for each $v\in I_{1}$; let $d_{1}$ and $d_{2}$ be two real numbers with $0<d_{1}<d_{2}$.
 Then $G$ admits a proper $\{-d_{1},0,d_{2}\}$-edge weighting.
 Proof.
-We would define an edge weighting such that each vertex in $I_{1}$ has weighted degree in $\{-3d_{1},d_{2}\}$ and each one in $I_{0}$ has weighted degree distinct from that in $\{-3d_{1},d_{2}\}$.
-In the beginning, every edge is weighted with 0. Let $I_{1}=\{x_{i}:i=1,2,\ldots,p\}$.
-First, label three of the edges incident to $x_{1}$ with $d_{1}$.
-Then $d_{w}(x_{1})=-3d_{1}$, and for each $y\in I_{0}$, $d_{w}(y)\in\{0,-d_{1}\}$; note that $\{0,-d_{1}\}$ is disjoint from $\{-3d_{1},d_{2}\}$.
+We construct an edge weighting $w$ such that every vertex in $I_{1}$ attains a weighted degree in $\{-3d_{1},d_{2}\}$ and each one in $I_{0}$ has weighted degree in $(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$.
+Initially, we assign weight 0 to every edge. Let $I_{1}=\{x_{i}:i=1,2,\ldots,p\}$.
+First, label three of the edges incident to $x_{1}$ with $d_{1}$. Then $d_{w}(x_{1})=-3d_{1}$, and for each $y\in I_{0}$, $d_{w}(y)\in\{0,-d_{1}\}\subseteq(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$.
 Second, for $i=2,3,\ldots,p$, suppose:
 (i) for each $j\geq i$, no edge incident to $x_{j}$ is weighted;
-(ii) for each vertex $y$ in $I_{0}$, $d_{w}(y)\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1},d_{2}\}$;
+(ii) for each vertex $y$ in $I_{0}$, $d_{w}(y)\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$;
 (iii) for each $j<i$, $d_{w}(x_{j})\in\{-3d_{1},d_{2}\}$.
 Case 1: If $x_{i}$ has a neighbor $y^{\ast}$ with $d_{w}(y^{\ast})\in(-\infty,-d_{1})\setminus\{-d_{2}-3d_{1}\}$, let $w(x_{i}y^{\ast})=d_{2}$.
-After this assignment, $d_{w}(x_{i})=d_{2}$, and the weighted degree of $y^{\ast}$ increases by $d_{2}$.
-Since before the assignment $d_{w}(y^{\ast})<-d_{1}$, after adding $d_{2}$ we have $d_{w}(y^{\ast})<d_{2}-d_{1}$.
-Moreover, because originally $d_{w}(y^{\ast})\neq-d_{2}-3d_{1}$, the new value satisfies $d_{w}(y^{\ast})\neq-3d_{1}$.
-Thus $d_{w}(y^{\ast})\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$, and one can verify that conditions (i), (ii), (iii) are maintained for $i+1$.
-Case 2: If each neighbor of $x_{i}$ has a weighted degree in $d_{1},d_{2}-d_{1})\cup\{-d_{2}-3d_{1}\}$, let $y_{1},y_{2},y_{3}$ be three of them.
-Let $w(x_{i}y_{1})=w(x_{i}y_{2})=w(x_{i}y_{3})=-d_{1}$.
-Then $d_{w}(x_{i})=-3d_{1}$ and each neighbor of $x_{i}$ has a weighted degree in $2d_{1},d_{2}-d_{1})\cup\{-d_{2}-3d_{1},-d_{2}-4d_{1}\}$ (this set is disjoint with $\{-3d_{1},d_{2}\}$), and one can check that (i), (ii), (iii) hold for $i+1$.
-Then finally, each vertex in $I_{0}$ has weighted degree in $(-\infty,d_{2}-d_{1})\setminus\{-3d_{1},d_{2}\}$, while each one in $I_{1}$ has weighted degree in $\{-3d_{1},d_{2}\}$, and we obtain a proper $\{-d_{1},0,d_{2}\}$-edge weighting.
+After this assignment, $d_{w}(x_{i})=d_{2}$, $d_{w}(y^{\ast})\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$, and each $y\in I_{0}\setminus\{y^{\ast}\}$ still satisfies $d_{w}(y)\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$.
+Thus, conditions (i)-(iii) are maintained for $i+1$.
+Case 2: If each neighbor of $x_{i}$ has a weighted degree in $d_{1},d_{2}-d_{1})\cup\{-d_{2}-3d_{1}\}$, let $y_{1},y_{2}$ and $y_{3}$ be three of them.
+Set $w(x_{i}y_{1})=w(x_{i}y_{2})=w(x_{i}y_{3})=-d_{1}$.
+Then $d_{w}(x_{i})=-3d_{1}$.
+For each $j=1,2,3$, $d_{w}(y_{j})\in[-2d_{1},d_{2}-2d_{1})\cup\{-d_{2}-4d_{1}\}$, which is a subset of $(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$.
+What is more, each $y\in I_{0}\setminus\{y_{1},y_{2},y_{3}\}$ still satisfies $d_{w}(y)\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$.
+Thus, conditions (i)-(iii) hold for $i+1$.
+Consequently, each vertex in $I_{0}$ attains a weighted degree in $(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$, while each one in $I_{1}$ has weighted degree in $\{-3d_{1},d_{2}\}$, and we obtain a proper $\{-d_{1},0,d_{2}\}$-edge weighting.
 ∎
@@
 Choose a maximum independent set $I_{0}$ in $G$.
-For $i\geq 1$, let $I_{i}$ be a maximum independent set in $G_{i}$, where $G_{i}=G[V\setminus(\cup_{j<i}I_{j})]$.
+For $i\geq 1$, let $I_{i}$ be a maximum independent set in $G_{i}$, where $G_{i}\triangleq G[V\setminus(\cup_{j<i}I_{j})]$.
 Let $\ell$ be the maximum integer such that $I_{\ell}\neq\emptyset$.
@@
 So suppose $2\leq\ell\leq k+1$.
-We will construct a proper $\{-d_{1},0,d_{2}\}$-edge weighting $w$ where construction proceeds in two phases. In the weighting process, for a vertex $v$ in $I_{i}$ ($i\in[1,\ell]$), we say $v$ is:
-  • Type I if $d_{w}(v)=(i-1)d_{2}$, and $v$ is incident to exactly (i1)(i-1) edges of weight $d_{2}$ connecting $v$ to $\cup_{j\geq 1}I_{j}$;
+We will construct a proper $\{-d_{1},0,d_{2}\}$-edge weighting $w$ in two phases. In the weighting process, for $i=1,2,\ldots,\ell,$ a vertex $v$ in $I_{i}$ is called:
+  • Type I if $d_{w}(v)=(i-1)d_{2}$, and $v$ is incident to exactly i1i-1 edges of weight $d_{2}$ connecting $v$ to $\cup_{j\geq 1}I_{j}$;
   • Type II if $d_{w}(v)=id_{2}-d_{1}$, and $v$ is incident to exactly $i$ edges of weight $d_{2}$ connecting $v$ to $\cup_{j\geq 1}I_{j}$ and one edge of weight $d_{1}$ connecting $v$ to $I_{0}$;
   • Type III if $v\in I_{i}$ for some $i\geq 2$, $d_{w}(v)=id_{2}-2d_{1}$, and $v$ is incident to exactly $i$ edges of weight $d_{2}$ connecting $v$ to $\cup_{j\geq 1}I_{j}$ and two edges of weight $d_{1}$: one connecting $v$ to $I_{0}$ and the other connecting $v$ to a vertex in $I_{1}$;
-  • hungry if $v\in I_{i}$ for some $i\geq 2$, $d_{w}(v)=jd_{2}$ for some $j\leq i-2$ and $v$ is incident to exactly $j$ edges of weight $d_{2}$ connecting $v$ to $\cup_{j\geq 1}I_{j}$;
-Goal. We will make sure $w$ satisfies the following conditions:
+  • hungry if $v\in I_{i}$ for some $i\geq 2$, $d_{w}(v)=pd_{2}$ for some $p\leq i-2$ and $v$ is incident to exactly $p$ edges of weight $d_{2}$ connecting $v$ to $\cup_{j\geq 1}I_{j}$.
+Goal. We will make sure $w$ finally satisfies the following conditions:
   • (1)
@@
   • (3)
-    for each $v$ in $I_{0}$, $d_{w}(v)<d_{2}-d_{1}$;
+    for each $y$ in $I_{0}$, $d_{w}(y)\in(-\infty,d_{2}-d_{1})$;
   • (4)
-    for each vertex $x$ in $I_{1}$ with $d_{w}(x)=-3d_{1}$, $x$ has at least two neighbors in $I_{0}$, and every neighbor $y$ of $x$ in $I_{0}$ satisfies $d_{w}(y)\neq-3d_{1}$;
+    for each vertex $x$ in $I_{1}$ with $d_{w}(x)=-3d_{1}$, $x$ has at least two neighbors in $I_{0}$, and every neighbor $y$ of $x$ in $I_{0}$ satisfies $d_{w}(y)\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$;
   • (5)
@@
   • (6)
-    for each vertex $x$ in $I_{1}$ with $d_{w}(x)=0$, every neighbor $y$ of $x$ in $I_{0}$ satisfies $d_{w}(y)\neq 0$;
+    for each vertex $x$ in $I_{1}$ with $d_{w}(x)=0$, every neighbor $y$ of $x$ in $I_{0}$ satisfies $d_{w}(y)\in(-\infty,d_{2}-d_{1})\setminus\{0\}$.
 Note that in (1), $\max\{(i-1)d_{2},id_{2}-2d_{1}\}<id_{2}-d_{1}$ because $d_{1}<d_{2}$.
 Moreover, for $1\leq i<j$, one has $id_{2}-d_{1}<\min\{(j-1)d_{2},jd_{2}-2d_{1}\}$.
-So (1), (2), (3) ensure there is no conflict in $G[\cup_{i\geq 1}I_{i}]$, and here is no conflict between $I_{0}$ and $\cup_{i\geq 2}I_{i}$.
-Clearly, (3), (4), (5), (6) ensure there is no conflict between $I_{0}$ and $I_{1}$.
-The construction proceeds in two phases. Phase 1 ensures condition (1), (2),(3) hold; Phase 2 ensures conditions (4), (5), (6) hold, with (1), (2),(3) remaining valid.
+So (1)-(3) ensure there is no conflict in $G[\cup_{i\geq 1}I_{i}]$, and there is no conflict between $I_{0}$ and $\cup_{i\geq 2}I_{i}$.
+Clearly, (3)-(6) ensure there is no conflict between $I_{0}$ and $I_{1}$.
+The construction proceeds in two phases. Phase 1 ensures condition (1)-(3) hold; Phase 2 ensures conditions (4)-(6) hold, with (1)-(3) remaining valid.
 In the beginning, every edge is weighted with 0.
@@
   • (2’)
-    for each $v$ in $I_{0}$, $d_{w}(v)\leq 0$;
+    for each $y$ in $I_{0}$, $d_{w}(y)\in(-\infty,0]$;
   • (3’)
     no Type I vertex in $I_{1}$ has a Type I neighbor in $\cup_{i\geq 2}I_{i}$.
+Clearly, the validity of conditions (1’)-(3’) guarantees that (1)-(3) hold.
 Step 1 By Lemma 2, there exists a matching $M_{\ell-1,\ell}$ in the bipartite graph $G[I_{\ell-1},I_{\ell}]$ that saturates $I_{l}$.
-Label each edge in $M_{\ell-1,\ell}$ with $d_{2}$.
+Weight each edge in $M_{\ell-1,\ell}$ with $d_{2}$.
 Then when $\ell=2$, each vertex in $I_{2}$ is Type I (has weighted degree $d_{2}=(2-1)d_{2}$); when $\ell\geq 3$, each vertex in $I_{\ell}$ is hungry (has weighted degree $d_{2}\leq(\ell-2)d_{2}$).
 Let $I^{\ast}_{\ell-1}$ be the set of vertices in $I_{\ell-1}$ that are saturated by $M_{\ell-1,\ell}$.
-Then the following holds.
+  • When $\ell=2$, by Lemma 2, there exists a matching $M^{\ast}_{0,1}$ in the bipartite graph $G[I_{0},I^{\ast}_{1}]$ that saturates $I^{\ast}_{1}$; weight each edge in $M^{\ast}_{0,1}$ with $d_{1}$, which makes each vertex in $I^{\ast}_{1}$ Type II (has weighted degree $d_{2}-d_{1}$), and each one in $I_{1}\setminus I^{\ast}_{1}$ Type I (has weighted degree $0=(1-1)d_{2}$).
+Then (1’) and (2’) already hold in this case.
   • When $\ell=3$, each vertex in $I^{\ast}_{2}$ is Type I (has weighted degree $d_{2}=(2-1)d_{2}$), each vertex in $I_{2}\setminus I^{\ast}_{2}$ is hungry (has weighted degree $0\leq(2-2)d_{2}$), and each (hungry) vertex $u$ in $I_{3}$ is incident to an edge of weight $d_{2}$ whose other endpoint lies in $I_{2}$.
   • When $\ell\geq 4$, each vertex in $I_{\ell-1}$ is hungry (has weighted degree at most $d_{2}\leq[(\ell-1)-2]d_{2}$), and each (hungry) vertex $u$ in $I_{\ell}$ is incident to an edge of weight $d_{2}$ whose other endpoint lies in $I_{\ell-1}$.
-  • When $\ell=2$, by Lemma 2, there exists a matching $M^{\ast}_{0,\ell-1}$ in the bipartite graph $G[I_{0},I^{\ast}_{1}]$ that saturates $I^{\ast}_{1}$; label each edge in $M^{\ast}_{0,1}$ with $d_{1}$, which makes each vertex in $I^{\ast}_{1}$ Type II (has weighted degree $d_{2}-d_{1}$), and each one in $I_{1}\setminus I^{\ast}_{1}$ Type I (has weighted degree $0=(1-1)d_{2}$).
-Then when $\ell=2$, one can check that (1’), (2’) already hold.
-Step 2 When $\ell\geq 3$, we process the indices $i=\ell-2,\ell-3,\ldots,1$ in decreasing order. At the start of step $i$, we assume:
+Step 2 When $\ell\geq 3$, we process the indices $i=\ell-2,\ell-3,\ldots,1$ in decreasing order. Right before proceeding $i$, we assume:
   • (A)
@@
   • (B)
-    every vertex in $I_{\alpha}$ for each $\alpha\geq i+1$ is Type I, Type II or hungry;
+    for each $\alpha\geq i+1$, every vertex in $I_{\alpha}$ is Type I, Type II or hungry;
   • (C)
-    for $\alpha\geq i+2$, if $I_{\alpha}$ has hungry vertices, then: for every hungry vertex $u$ in $I_{\alpha}$ and every $\beta\in[i+1,\alpha-1]$, $u$ is incident to at least one edge of weight $d_{2}$ whose other endpoint lies in $I_{\beta}$.
+    for $\alpha\geq i+2$, if $I_{\alpha}$ has hungry vertices, then: for every hungry vertex $u$ in $I_{\alpha}$ and every $\beta=i+1,\ldots,\alpha-1$, $u$ is incident to at least one edge of weight $d_{2}$ whose other endpoint lies in $I_{\beta}$.
 (The starting point of this inductive hypothesis holds true after Step 1.)
@@
 After processing all edges in $E_{i+1}$, let $U_{i+1}$ be the set of hungry vertices in $\cup_{j\geq i+1}I_{j}$.
-By construction, if an edge has both endpoints in $U_{i+1}$, then it has weight $d_{2}$; hence the induced subgraph $G[U_{i+1}]$ (if nonempty) has all edges weighted with $d_{2}$.
-Take an arbitrary vertex $u\in U_{i+1}$, and suppose $u\in I_{\alpha}$ for some $\alpha\in[i+1,\ell]$.
+By construction, if an edge has both endpoints in $U_{i+1}$, then it has already been weighted with $d_{2}$; hence the induced subgraph $G[U_{i+1}]$ (if nonempty) has all edges weighted with $d_{2}$.
+Take an arbitrary vertex $u\in U_{i+1}$, and suppose $u\in I_{\alpha}$ for some $\alpha=i+1,\ldots,\ell$.
 Since $u$ is hungry, $d_{w}(u)=jd_{2}$ for some $j\leq\alpha-2$.
-By assumption (C), for each $\beta\in[i+1,\alpha-1]$, $u$ is incident to at least one edge of weight $d_{2}$ with other endpoint in $I_{\beta}$.
+By assumption (C), for each $\beta=i+1,\ldots,\alpha-1$, $u$ is incident to at least one edge of weight $d_{2}$ with other endpoint in $I_{\beta}$.
 There are $(\alpha-1)-(i+1)+1=\alpha-i-1$ such indices $\beta$.
-Thus at most $(\alpha-2)-(\alpha-i-1)=i-1$ edges of weight $d_{2}$ incident to $u$ can have their other endpoint in $\cup_{j\geq\alpha+1}I_{j}$.
+Thus at most $(\alpha-2)-(\alpha-i-1)=i-1$ edges of weight $d_{2}$ incident to $u$ can have their other endpoints in $\cup_{j\geq\alpha+1}I_{j}$.
 Recall that $G[U_{i+1}]$ has all edges weighted with $d_{2}$.
@@
 Each such independent set is also independent in $G$.
-Hence we can write $U_{i+1}=J_{1}\cup\cdots\cup J_{r}$ with $r\leq i$, where each $J_{t}$ is independent in $G$.
-By Lemma 2, for each $t\in[1,r]$ there exists a matching $M^{\prime}_{i,t}$ in the bipartite graph $G[I_{i},J_{t}]$ that saturates $J_{t}$.
+Hence we can write $U_{i+1}=J_{1}\oplus\cdots\oplus J_{r}$ with $r\leq i$, where each $J_{t}$ is independent in $G$.
+By Lemma 2, for each $t=1,2,\ldots,r$, there exists a matching $M^{\prime}_{i,t}$ in the bipartite graph $G[I_{i},J_{t}]$ that saturates $J_{t}$.
 For every edge $e$ in $\bigcup_{t=1}^{r}M^{\prime}_{i,t}$, set $w(e)=d_{2}$.
-At this point, each vertex $v\in I_{i}$ is incident to at most $i$ edges of weight $d_{2}$ from these matchings, so $d_{w}(v)=jd_{2}$ for some $j\leq r\leq i$.
+At this point, each vertex $v\in I_{i}$ is incident to at most $r$ edges of weight $d_{2}$ from these matchings, so $d_{w}(v)=jd_{2}$ for some $j\leq r\leq i$.
 Let $\overline{I}_{i}\subseteq I_{i}$ be the set of vertices with weighted degree exactly $id_{2}$.
@@
   • (B’)
-    every vertex in $I_{\alpha}$ for each $\alpha\geq i$ is Type I, Type II or hungry;
+    for each $\alpha\geq i$, every vertex in $I_{\alpha}$ is Type I, Type II or hungry;
   • (C’)
-    for $\alpha\geq i+1$, if $I_{\alpha}$ has hungry vertices, then: for every hungry vertex $u$ in $I_{\alpha}$ and every $\beta\in[i,\alpha-1]$, $u$ is incident to at least one edge of weight $d_{2}$ whose other endpoint lies in $I_{\beta}$.
+    for $\alpha\geq i+1$, if $I_{\alpha}$ has hungry vertices, then: for every hungry vertex $u$ in $I_{\alpha}$ and every $\beta=i,\ldots,\alpha-1$, $u$ is incident to at least one edge of weight $d_{2}$ whose other endpoint lies in $I_{\beta}$.
 Indeed, (B’) holds because: if $v\in\cup_{j\geq i+1}I_{j}$ was already Type I or II after step $i+1$, it remains Type I or II; if $u\in U_{i+1}$ received an edge of weight $d_{2}$ from some $M^{\prime}_{i,t}$, it now becomes Type I or remains hungry (the extra $d_{1}$ edges to $I_{0}$ do not affect the condition for vertices in $U_{i+1}$); if $v\in I_{i}$, then either it received at most i1i-1 edges of weight $d_{2}$ from matchings $M^{\prime}_{i,t}$’s (in which case it becomes Type I or remains hungry), or received exactly $i$ edges of weight $d_{2}$ and exactly one edge of weight $d_{1}$ from $\overline{M}_{0,i}$ (in which case it becomes Type II).
-Proceeding in this way for $i=\ell-2,\ell-3,\ldots,1$, we obtain at the end of Phase 1 that every vertex in $\cup_{\alpha\geq 1}I_{\alpha}$ is Type I or II.
+Proceeding in this way for $i=\ell-2,\ell-3,\ldots,1$, we obtain at the end that every vertex in $\cup_{\alpha\geq 1}I_{\alpha}$ is Type I or II.
 Since at the end, on one hand, each $v$ in $I_{1}$ has weighted degree $d_{2}-d_{1}$ (Type II) or 0 (Type I).
-On the other hand, if some $u$ in $I_{\alpha}$ ($\alpha\geq 2$) is hungry, it should satisfies (C’): $u$ is incident to at least one edge of weight $d_{2}$ with other endpoint in $I_{\beta}$ for each $\beta\in[1,\alpha-1]$ (at the end $i=1$).
-That is, $d_{w}(u)\geq(\alpha-1)d_{2}$ and $u$ is not hungry, a contradiction.
-Consequently, condition (1’) holds and there is no conflict within the subgraph $G[\cup_{i\geq 1}I_{i}]$.
-Moreover, every vertex $v\in I_{0}$ satisfies $d_{w}(v)\leq 0$ at this stage and (2’) holds.
-For (3’), if there exists a Type I vertex $v$ in $I_{1}$ that has a Type I neighbor $u$ in some $I_{i}$ ($i\geq 2$), let $e_{v}$ and $e_{u}$ be the edges joining $v$ and $u$, respectively, to $I_{0}$ (recall that each vertex in $\cup_{j\geq 1}I_{j}$ has a neighbor in $I_{0}$ since $I_{0}$ is a maximum independent set).
-Clearly, $e_{v},e_{u},uv$ have not been weighted since $v,u$ are Type I (moreover, $v$ is Type I and in $I_{1}$).
+On the other hand, if some $u$ in $I_{\alpha}$ ($\alpha\geq 2$) is hungry, it should satisfies (C’): $u$ is incident to at least one edge of weight $d_{2}$ with other endpoint in $I_{\beta}$ for each $\beta=1,\ldots,\alpha-1$ (at the end $i=1$).
+That is, $d_{w}(u)\geq(\alpha-1)d_{2}$ which implies that $u$ is not hungry, a contradiction.
+Consequently, condition (1’) holds.
+Moreover, every $y$ in $I_{0}$ satisfies $d_{w}(y)\in(-\infty,0]$ at this stage and (2’) holds.
+For (3’), if there exists a Type I vertex $v$ in $I_{1}$ that has a Type I neighbor $u$ in some $I_{i}$ ($i\geq 2$), let $e_{v}$ and $e_{u}$ be the edges joining $v$ and $u$ to $I_{0}$, respectively (recall that each vertex in $\cup_{j\geq 1}I_{j}$ has a neighbor in $I_{0}$ since $I_{0}$ is a maximum independent set).
+Clearly, $e_{v},e_{u}$ and $uv$ have not been weighted since $v$ and $u$ are both Type I (moreover, $v$ is in $I_{1}$).
 Set $w(e_{v})=w(e_{u})=-d_{1}$ and set $w(uv)=d_{2}$.
-Then both $u,v$ become Type II (so condition (1’) remains valid), and every vertex in $I_{0}$ still has weighted degree at most 0 (so condition (2’) remains valid).
-Repeat this step until no Type I vertex in $I_{1}$ has a Type I neighbor in $\cup_{i\geq 2}I_{i}$, and (3’) holds.
+Then both $u$ and $v$ become Type II (so condition (1’) remains valid), and every vertex in $I_{0}$ still has weighted degree at most 0 (so condition (2’) remains valid).
+Repeat this step until no Type I vertex in $I_{1}$ has a Type I neighbor in $\cup_{i\geq 2}I_{i}$, and (3’) holds finally.
 Phase 2: Conflict resolution
-For $i=0,1$, let $I_{i,0}=\{v\in I_{i}:d_{w}(v)=0\}.$ After Phase 1, by (1’) and (2’), the only possible conflicts in $G$ are those belonging to the bipartite subgraph $G[I_{0,0},I_{1,0}]$.
+For $i=0,1$, let $I_{i,0}=\{v\in I_{i}:d_{w}(v)=0\}.$ By our construction, all edges incident to vertices in $I_{0,0}\cup I_{1,0}$ currently have weight 0.
+After Phase 1, by (1’) and (2’), the only possible conflicts in $G$ are those belonging to the bipartite subgraph $G[I_{0,0},I_{1,0}]$.
 If $G[I_{0,0},I_{1,0}]$ is empty, we are done.
 Assume therefore that it is nonempty.
-For each vertex $x\in I_{1,0}$, choose a neighbor $y_{x}$ (if exists) of $x$ in $I_{0,0}$.
-Remark on dynamics: During Phase 2, the sets $I_{1,0}$, $I_{0,0}$, and the classifications based on them will change as we assign new weights.
-We will carefully track these changes to ensure the algorithm terminates.
-At this point, every vertex in $I_{1,0}$ has all its neighbors either in $I_{0}$ or in Type II vertices of $\cup_{i\geq 2}I_{i}$ by (1’) and (3’).
-We classify the vertices of $I_{1,0}$ according to their number of neighbors in $I_{0}$:
-$I_{1,0}=X_{2^{+}}\cup X_{1}=X_{3^{+}}\cup X_{2}\cup X_{1},$
-where $X_{i^{+}}=\{x\in I_{1,0}:|N(x)\cap I_{0}|\geq i\}$ and $X_{i}=\{x\in I_{1,0}:|N(x)\cap I_{0}|=i\}$ for $i\geq 1$.
-Now consider the bipartite subgraph $H^{\prime}:=G[X_{2}\cup X_{1},\cup_{i\geq 2}I_{i}]$. On one hand, for any $x\in X_{2}\cup X_{1}$:
-  • If $x\in X_{2}$, then $|N(x)\cap I_{0}|=2$, so $x$ has at least k2k-2 neighbors in $\cup_{i\geq 2}I_{i}$ (since all neighbors are either in $I_{0}$ or Type II vertices of $\cup_{i\geq 2}I_{i}$).
-  • If $x\in X_{1}$, then $|N(x)\cap I_{0}|=1$, so $x$ has at least $k-1\geq k-2$ neighbors in $\cup_{i\geq 2}I_{i}$.
-For any Type II vertex $z\in\cup_{i\geq 2}I_{i}$, by definition it is incident to exactly one edge of weight $d_{1}$ (to a vertex in $I_{0}$) and exactly $i$ edges of weight $d_{2}$ for some $i\geq 2$.
-Since all edges incident to vertices in $I_{1,0}$ currently have weight 0, any edge of weight $d_{2}$ incident to $z$ must connect $z$ to a vertex not in $I_{1,0}$.
-Thus $z$ has at least one neighbor not in $I_{1,0}$ via an edge of weight $d_{2}$, plus the edge of weight $d_{1}$ to $I_{0}$.
-Consequently, $z$ can be adjacent to at most k2k-2 vertices in $I_{1,0}$ (so to at most k2k-2 vertices in $X_{2}\cup X_{1}$).
-Thus, in $H^{\prime}$, every vertex in $X_{2}\cup X_{1}$ has degree at least k2k-2, and every vertex in $\cup_{i\geq 2}I_{i}$ has degree at most k2k-2.
-To verify Hall’s condition for a matching saturating $X_{2}\cup X_{1}$, take any subset $S\subseteq X_{2}\cup X_{1}$.
-The number of edges incident to $S$ in $H^{\prime}$ is at least |S|(k2)|S|(k-2).
-Since each vertex in the neighborhood $N_{H^{\prime}}(S)$ has degree at most k2k-2, we have
-$|S|(k-2)\leq e(S,N_{H^{\prime}}(S))\leq|N_{H^{\prime}}(S)|(k-2).$
-Dividing by k2>0k-2>0 (as $k\geq 3$) gives $|S|\leq|N_{H^{\prime}}(S)|$.
-Hence Hall’s condition holds, and there exists a matching $M$ in $H^{\prime}$ that saturates $X_{2}\cup X_{1}$.
-Write $M=M_{1}\cup M_{2}$, where $M_{i}$ saturates $X_{i}$ for $i=1,2$.
+For each $x\in I_{1,0}$ with $N(x)\cap I_{0,0}\neq\emptyset$, choose an arbitrary vertex from $N(x)\cap I_{0}$ and denote it by $y_{x}$.
+For $i\geq 1$, let $X_{i}=\{x\in I_{1,0}:|N(x)\cap I_{0}|=i\}$ and $X_{i^{+}}=\cup_{j\geq i}X_{j}$.
+Now consider the bipartite subgraph $H^{\prime}:=G[X_{1}\cup X_{2},\cup_{i\geq 2}I_{i}]$.
+On one hand, for any $x\in X_{1}\cup X_{2}$, the vertex $x$ has at least k2k-2 Type II neighbors in $\bigcup_{i\geq 2}I_{i}$.
+This holds because, by conditions (1’) and (3’), every neighbor of $x$ either lies in $I_{0}$ or is a Type II vertex in $\bigcup_{i\geq 2}I_{i}$.
+On the other hand, consider a Type II vertex $z\in I_{i}$ with $i\geq 2$.
+By definition, $z$ is incident to exactly $i$ edges of weight $d_{2}$ and one edge of weight $d_{1}$.
+Since all edges incident to vertices in $X_{1}\cup X_{2}$ currently carry weight 0, any edge of weight $d_{2}$ or $d_{1}$ incident to $z$ must have its other endpoint outside $X_{1}\cup X_{2}$.
+Thus $z$ can be adjacent to at most k2k-2 vertices in $X_{1}\cup X_{2}$.
+Thus, in $H^{\prime}$, every vertex in $X_{1}\cup X_{2}$ has degree at least k2k-2, and every vertex in $\cup_{i\geq 2}I_{i}$ has degree at most k3k-3.
+By Hall’s theorem, there exists a matching $M$ in $H^{\prime}$ that saturates $X_{1}\cup X_{2}$.
+Write $M=M_{1}\oplus M_{2}$, where $M_{i}$ saturates $X_{i}$ for $i=1,2$.
+For each $x\in X_{1}\cup X_{2}$, let $z_{x}$ be the (Type II) vertex in $\cup_{i\geq 2}I_{i}$ such that $xz_{x}\in M$.
 We will resolve the conflicts involving $X_{2^{+}}$ and $X_{1}$ separately.
 1. Resolving conflicts incident to $X_{2^{+}}$.
-Let $X^{\ast}_{2^{+}}\subseteq X_{2^{+}}$ be the set of vertices that have a neighbor in $I_{0,0}$. If $X^{\ast}_{2^{+}}=\emptyset$, then we are done.
-We process vertices from $X_{2^{+}}$ iteratively.
-Each iteration removes one vertex from $X_{2^{+}}$ (and hence from $X^{\ast}_{2^{+}}$).
-Although $X^{\ast}_{2^{+}}$ may expand due to vertices re-entering $I_{0,0}$ in Case 1, the set $X_{2^{+}}$ itself strictly decreases in size with each iteration.
-Therefore, the loop terminates after at most $|X_{2^{+}}|$ iterations.
-Let $i=1$.
-While $X^{\ast}_{2^{+}}\neq\emptyset$, we select a vertex $x_{i}$ from $X^{\ast}_{2^{+}}$ (arbitrarily for $i=1$, and from the updated $X^{\ast}_{2^{+}}$ for $i>1$) and handle it according to the cases below.
-After processing $x_{i}$, it is removed from $I_{1,0}$ (since its weighted degree becomes non-zero), and consequently from $X_{2^{+}}$ and $X^{\ast}_{2^{+}}$ (this process will gradually eliminates vertices from $X_{2^{+}}$ until $X^{\ast}_{2^{+}}$ becomes empty).
+Let $X^{\ast}_{2^{+}}\subseteq X_{2^{+}}$ be the set of vertices that have a neighbor in $I_{0,0}$. If $X^{\ast}_{2^{+}}=\emptyset$, then we are done. Suppose $X^{\ast}_{2^{+}}\neq\emptyset$.
+We select a vertex $x_{i}$ from $X^{\ast}_{2^{+}}$ (arbitrarily for $i=1$, and from the updated $X^{\ast}_{2^{+}}$ for $i>1$) and handle it according to the cases below.
+After processing $x_{i}$, it will be removed from $I_{1,0}$ (since its weighted degree will become non-zero), and consequently from $X_{2^{+}}$ and $X^{\ast}_{2^{+}}$ (this process will gradually eliminates vertices from $X_{2^{+}}$ until $X^{\ast}_{2^{+}}$ becomes empty).
 We then update $I_{1,0}$, $X_{2^{+}}$, $X^{\ast}_{2^{+}}$ accordingly.
@@
   • (ii)
-    for every $y\in I_{0}$, $d_{w}(y)<d_{2}-d_{1}$;
+    for every $y\in I_{0}$, $d_{w}(y)\in(-\infty,d_{2}-d_{1})$;
   • (iii)
-    for each previously processed vertex $x_{j}$ ($j<i$), one of the following items holds:
-    • (i’)
+    each previously processed vertex $x_{j}$ ($j<i$) satisfies one of the following conditions:
     • $d_{w}(x_{j})=d_{2}-d_{1}$;
-    • (ii’)
-      $d_{w}(x_{j})=-3d_{1}$ and every neighbor $y$ of $x_{j}$ in $I_{0}$ satisfies $d_{w}(y)\neq-3d_{1}$.
+    • $d_{w}(x_{j})=-3d_{1}$ and every neighbor $y$ of $x_{j}$ in $I_{0}$ satisfies $d_{w}(y)\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$.
 (For $i=1$, conditions (i) and (ii) hold; no vertex has been processed and so condition (iii) holds vacuously.)
-We distinguish three cases.
-Case 1: $x_{i}\in X_{2^{+}}$, and $x_{i}$ has a neighbor $y^{\ast}\in I_{0}$ with $d_{w}(y^{\ast})\in(-\infty,-d_{1})\setminus\{-d_{2}-3d_{1}\}$.
+We distinguish two cases.
+Case 1: If there exists $y^{\ast}\in N(x_{i})\cap I_{0}$ such that $d_{w}(y^{\ast})\in(-\infty,-d_{1})\setminus\{-d_{2}-3d_{1}\}$.
+Clearly, $y^{\ast}\neq y_{x_{i}}$, since by the definition of $y_{x_{i}}$, we have $d_{w}(y_{x_{i}})=0$ at this stage.
 Set $w(x_{i}y^{\ast})=d_{2}$ and $w(x_{i}y_{x_{i}})=-d_{1}$.
-After these assignments, $d_{w}(x_{i})=d_{2}-d_{1}$ (hence condition (iii)(i’) holds for $x_{i}$), $d_{w}(y_{x_{i}})=-d_{1}$, and the weighted degree of $y^{\ast}$ increases by $d_{2}$.
-Since before the assignment $d_{w}(y^{\ast})<-d_{1}$, after adding $d_{2}$ we have $d_{w}(y^{\ast})<d_{2}-d_{1}$, which fulfills condition (ii).
-Furthermore, because originally $d_{w}(y^{\ast})\neq-d_{2}-3d_{1}$, we have $d_{w}(y^{\ast})\neq-3d_{1}$.
-Thus $d_{w}(y^{\ast})\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$, and conditions (ii) and (iii) continue to hold for $i+1$.
-Case 2: $x_{i}\in X_{3^{+}}$ and every neighbor of $x_{i}$ in $I_{0}$ has weighted degree in $d_{1},d_{2}-d_{1})\cup\{-d_{2}-3d_{1}\}$.
-Since $x_{i}\in X_{3^{+}}$, it has at least three neighbors in $I_{0}$.
-Arbitrarily choose three edges $x_{i}y_{1},x_{i}y_{2},x_{i}y_{3}$ in $G[I_{0},X_{3^{+}}]$ and set $w(x_{i}y_{1})=w(x_{i}y_{2})=w(x_{i}y_{3})=-d_{1}$.
+After these assignments, $d_{w}(x_{i})=d_{2}-d_{1}$, while $d_{w}(y_{x_{i}})=-d_{1}$ and $d_{w}(y^{\ast})\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$.
+Moreover, the weighted degree of each vertex in $(I_{0}\setminus\{y_{x_{i}},y^{\ast}\})\cup(X_{2^{+}}\setminus\{x_{i}\})$ remains unchanged.
+Thus, conditions (i)-(iii) continue to hold for $i+1$.
+Case 2: If every vertex in $N(x_{i})\cap I_{0}$ has weighted degree in $d_{1},d_{2}-d_{1})\cup\{-d_{2}-3d_{1}\}$.
+Subcase 2.1: If $x_{i}\in X_{3^{+}}$.
+Let $y_{1},y_{2},y_{3}\in N(x_{i})\cap I_{0}$ and set $w(x_{i}y_{1})=w(x_{i}y_{2})=w(x_{i}y_{3})=-d_{1}$.
 This gives $d_{w}(x_{i})=-3d_{1}$, while $d_{w}(y_{1}),d_{w}(y_{2}),d_{w}(y_{3})\in[-2d_{1},d_{2}-2d_{1})\cup\{-d_{2}-4d_{1}\}$.
-Then every neighbor of $x_{i}$ in $I_{0}$ has weighted degree in $2d_{1},d_{2}-d_{1})\cup\{-d_{2}-4d_{1},-d_{2}-3d_{1}\}$; this set is consistent with (ii), and does not include $3d_{1}$ and so (iii) also holds for $i+1$.
-Case 3: $x_{i}\in X_{2}$ and every neighbor of $x_{i}$ in $I_{0}$ has weighted degree in $d_{1},d_{2}-d_{1})\cup\{-d_{2}-3d_{1}\}$.
-Let $y^{\ast}$ be the unique neighbor of $x_{i}$ in $I_{0}$ other than $y_{x_{i}}$.
-Let $x_{i}z$ be the edge in $M_{2}$ that saturates $x_{i}$, where $z\in I_{q}$ for some $q\geq 2$.
-By (1’) and (3’), $z$ is Type II and so $d_{w}(z)=qd_{2}-d_{1}$.
-Now set $w(x_{i}y_{x_{i}})=w(x_{i}y^{\ast})=w(x_{i}z)=-d_{1}$.
-This gives $d_{w}(x_{i})=-3d_{1}$, while $d_{w}(y_{x_{i}})=-d_{1}$, $d_{w}(z)=qd_{2}-2d_{1}$ (that is, $z$ becomes Type III and so condition (1) remains valid), and $d_{w}(y^{\ast})\in[-2d_{1},d_{2}-2d_{1})\cup\{-d_{2}-4d_{1}\}$.
-Then $\{d_{w}(y^{\ast}),d_{w}(y_{x_{i}})\}\subseteq[-2d_{1},d_{2}-d_{1})\cup\{-d_{2}-4d_{1}\}$; this set is consistent with (ii), and does not include $3d_{1}$ and so (iii) also holds for $i+1$.
+Then every vertex in $N(x_{i})\cap I_{0}$ has weighted degree in $2d_{1},d_{2}-d_{1})\cup\{-d_{2}-4d_{1},-d_{2}-3d_{1}\}$, which is a subset of $(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$.
+Moreover, the weighted degree of each vertex in $(I_{0}\setminus\{y_{1},y_{2},y_{3}\})\cup(X_{2^{+}}\setminus\{x_{i}\})$ remains unchanged.
+Thus, conditions (i)-(iii) continue to hold for $i+1$.
+Subcase 2.2: If $x_{i}\in X_{2}$.
+Assume $N(x_{i})\cap I_{0}=\{y_{x_{i}},y^{\prime}\}$ and $z_{x_{i}}\in I_{q}$ for some $q\geq 2$.
+Recall that $z_{x_{i}}$ is Type II and so $d_{w}(z_{x_{i}})=qd_{2}-d_{1}$.
+Now set $w(x_{i}y_{x_{i}})=w(x_{i}y^{\prime})=w(x_{i}z_{x_{i}})=-d_{1}$.
+This gives $d_{w}(x_{i})=-3d_{1}$, while $d_{w}(y_{x_{i}})=-d_{1}$, $d_{w}(y^{\prime})\in[-2d_{1},d_{2}-2d_{1})\cup\{-d_{2}-4d_{1}\}$ and $d_{w}(z_{x_{i}})=qd_{2}-2d_{1}$ (that is, $z_{x_{i}}$ becomes Type III and so condition (1) remains valid).
+Then $\{d_{w}(y_{x_{i}}),d_{w}(y^{\prime})\}\subseteq[-2d_{1},d_{2}-2d_{1})\cup\{-d_{2}-4d_{1}\}$, which is a subset of $(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$.
+Moreover, the weighted degree of each vertex in $(I_{0}\setminus\{y_{x_{i}},y^{\prime}\})\cup(X_{2^{+}}\setminus\{x_{i}\})$ remains unchanged.
+Thus, conditions (i)-(iii) continue to hold for $i+1$.
 Remark on the dynamics of $I_{0,0}$: In Case 1, it is possible that $d_{w}(y^{\ast})$ becomes 0 after adding $d_{2}$ (if originally $d_{w}(y^{\ast})=-d_{2}$), placing $y^{\ast}$ back into $I_{0,0}$.
 However, by assumption (i), all edges incident to $\{x_{j}:j>i\}$ are still unweighted, so any new conflict involving $y^{\ast}$ and $x_{j}$ ($j>i$) will be resolved when $x_{j}$ is processed.
-Moreover, $y^{\ast}$ cannot conflict with already processed vertices $x_{p}$ ($p<i$) because those have weighted degrees either $d_{2}-d_{1}$ or $3d_{1}$, both different from 0.
-For cases 2 and 3, all modified weighted degrees in $I_{0}$ become negative; so $I_{0}$-vertices cannot re-enter $I_{0,0}$.
+Moreover, $y^{\ast}$ cannot conflict with already processed vertices $x_{s}$ ($s<i$) because those have weighted degrees either $d_{2}-d_{1}$ or $3d_{1}$, both different from 0.
+For Case 2, all modified weighted degrees in $I_{0}$ become negative; so $I_{0}$-vertices cannot re-enter $I_{0,0}$.
+After handling $x_{i}$, update $I_{0,0},X_{2^{+}}$ and $X^{\ast}_{2^{+}}$, respectively.
+Note that $x_{i}$ and all previously processed vertices $x_{j}$ ($j<i$) are no longer in $X^{\ast}_{2^{+}}$ (no longer in $X_{2^{+}}$ either), since they either have weighted degree $d_{2}-d_{1}$ (which differs from 0), or have weighted degree $3d_{1}$ while all their neighbors in $I_{0}$ have weighted degrees different from $3d_{1}$.
+If $X^{\ast}_{2^{+}}=\emptyset$, we are done.
+If $X^{\ast}_{2^{+}}\neq\emptyset$ let $x_{i+1}\in X^{\ast}_{2^{+}}$ and the assumptions (i)-(iii) hold for $i+1$.
+Proceeding and update $X^{\ast}_{2^{+}}$ repeatedly, we eventually eliminate all conflicts in $G[I_{0},X_{2^{+}}]$.
+Moreover, conditions (1)-(3) remain valid currently.
+Hence no conflict remains incident to $X_{2^{+}}$ finally.
 Remark on the dynamics of $X^{\ast}_{2^{+}}$: Since (in Case 1) $I_{0,0}$ may expand, $X^{\ast}_{2^{+}}$ may also expand.
 However, $X^{\ast}_{2^{+}}$ is included in $X_{2^{+}}$, and $X_{2^{+}}$ decreases at each step.
-So the recursive process will terminate with $X_{2^{+}}=\emptyset$.
-After handling $x_{i}$, update $I_{0,0},X_{2^{+}}$ and $X^{\ast}_{2^{+}}$ respectively.
-Note that $x_{i}$ and all previously processed vertices $x_{p}$ ($p<i$) are no longer in $X^{\ast}_{2^{+}}$ (no longer in $X_{2^{+}}$ either), since they either have weighted degree $d_{2}-d_{1}$ (which differs from 0), or have weighted degree $3d_{1}$ while all their neighbors in $I_{0}$ have weighted degrees different from $3d_{1}$.
-If $X^{\ast}_{2^{+}}=\emptyset$, we are done.
-If $X^{\ast}_{2^{+}}\neq\emptyset$ let $x_{i+1}\in X^{\ast}_{2^{+}}$ and the assumptions (i)–(iii) hold for $i+1$.
-Proceeding and update $X^{\ast}_{2^{+}}$ repeatedly, we eventually eliminate all conflicts in $G[I_{0},X_{2^{+}}]$.
-Moreover, (1), (2), (3) remain valid currently.
-Hence no conflict remains incident to $X_{2^{+}}$.
+So the recursive process will terminate with $X^{\ast}_{2^{+}}=\emptyset$.
 2. Resolving conflicts incident to $X_{1}$.
@@
 Note that $X_{1}$ remains unchanged.
-Let $X^{\ast}_{1}\subseteq X_{1}$ be the set of vertices in the current $X_{1}$ that have a neighbor in (updated) $I_{0,0}$.
+Let $X^{\ast}_{1}\subseteq X_{1}$ be the set of vertices in $X_{1}$ that have a neighbor in (updated) $I_{0,0}$.
 If $X^{\ast}_{1}=\emptyset$, we are finished.
 Assume $X^{\ast}_{1}\neq\emptyset$ and let $M^{\ast}_{1}$ be the subset of $M_{1}$ consisting of edges incident to vertices in $X^{\ast}_{1}$.
-Recall that each edge in $M^{\ast}_{1}$ connects a vertex $x\in X^{\ast}_{1}$ to a Type II vertex $z\in\cup_{i\geq 2}I_{i}$, and each edge in $M^{\ast}_{1}$ is unweighted.
 Assign weight $d_{1}$ to every edge in $M^{\ast}_{1}$.
-After the assignment:
-  • $d_{w}(x)$ changes from 0 to $d_{1}$;
-  • $d_{w}(y_{x})$ remains 0 (since edge $xy_{x}$ is not in $M^{\ast}_{1}$ and keeps weight 0);
-  • $d_{w}(z)$ decreases by $d_{1}$ for the corresponding $z\in\cup_{i\geq 2}I_{i}$ which becomes Type III.
-This implies that condition (5) holds, and (1), (2), (3) remain valid.
-Note that, at present, both $X^{\ast}_{2^{+}}$ and $X^{\ast}_{1}$ are empty, and so (6) holds.
-Condition (4) holds because: vertices with $d_{w}(x)=-3d_{1}$ come from Case 2 or Case 3.
-In Case 2, $x\in X_{3^{+}}$ has at least three neighbors in $I_{0}$; in Case 3, $x\in X_{2}$ has exactly two neighbors in $I_{0}$.
+After the assignment, for each $x\in X^{\ast}_{1}$, $d_{w}(x)=-d_{1}$, while $d_{w}(y_{x})$ remains 0 and $d_{w}(z_{x})$ decreases by $d_{1}$, making $z_{x}$ a Type III vertex.
+This implies that condition (5) holds, and conditions (1)-(3) remain valid.
+Note that, at present, both $X^{\ast}_{2^{+}}$ and $X^{\ast}_{1}$ are empty, and so condition (6) holds.
+Condition (4) holds because: vertices with $d_{w}(x)=-3d_{1}$ come from Case 2.
+In Case 2.1, $x\in X_{3^{+}}$ has at least three neighbors in $I_{0}$; in Case 2.2, $x\in X_{2}$ has exactly two neighbors in $I_{0}$.
 In both cases, all neighbors $y$ of $x$ in $I_{0}$ satisfy $d_{w}(y)\neq-3d_{1}$ by construction.
-Thus, $w$ is a proper $\{-d_{1},0,d_{2}\}$-edge weighting of $G$.
-This completes the proof of Theorem 1.
-∎
+Thus, conditions (1)-(6) all hold, and $w$ is a proper $\{-d_{1},0,d_{2}\}$-edge weighting of $G$.
+This completes the proof of Theorem 1. ∎
 3. Concluding Remarks
-Theorem 1 and Corollary 2 establish that every regular graph is universally 3-edge-weightable.
-The independent-set decomposition technique, originally developed for total weightings, proves adaptable to edge-weightings with arbitrary 3-element set of weights.
-Our work suggests that regularity provides a structural framework within which arbitrary three-element weight set can be accommodated.
-The conflict-resolution scheme developed here may find applications in other weighting and labeling problems.
-The two-phase approach—global matching adjustments followed by local conflict resolution—offers a template for handling parameter-dependent constructions in combinatorial graph theory.
-Future work. Several natural questions arise from our result:
+This paper establishes that every regular graph is universally 3-edge-weightable.
+The conflict-resolution scheme introduced in this work may be applicable to other weighting and labeling problems.
+The two-phase approach–comprising global matching adjustments followed by local conflict resolution–provides a framework for handling parameter-dependent constructions in combinatorial graph theory.
+Future work. Our result naturally leads to several open questions:
   • Does universal 3-edge-weightability extend to all nice graphs? (Problem 1.1)